### v2

# Fast Crystal Tensor Property Prediction: A General O(3)-Equivariant Framework Based on Polar Decomposition

Abstract
Predicting the tensor properties of crystalline materials is a fundamental task in materials science. Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance. This equivariance constraint often introduces tremendous computational costs, necessitating specialized designs for effective and efficient predictions. To address this limitation, we propose a general $O(3)$-equivariant framework for fast crystal tensor property prediction, called GoeCTP. Our framework is efficient as it does not need to impose equivalence constraints onto the network architecture. Instead, GoeCTP captures the tensor equivariance with a simple external rotation and reflection (R&R) module based on polar decomposition. The crafted external R&R module can rotate and reflect the crystal into an invariant standardized crystal position in space without introducing extra computational cost. We show that GoeCTP is general as it is a plug-and-play module that can be smoothly integrated with any existing single-value property prediction framework for predicting tensor properties. Experimental results indicate that GoeCTP achieves higher prediction performance and runs 13$\times$ faster compared to existing state-of-the-art methods in elastic benchmarking datasets, underscoring its effectiveness and efficiency.
1 Introduction
The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering (Yan et al., 2024b). Compared to single-value property prediction, predicting tensor properties of crystalline materials is substantially more complex. This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress (Nye, 1985; Resta, 1994; Yan et al., 2024b). Consequently, tensor property prediction modeling necessitates preserving consistency with the crystal’s spatial position, exhibiting “special” $O(3)$ equivariance (Yan et al., 2024b). Furthermore, high-order tensor property prediction is computationally intensive due to its high dimensionality and large data volume. Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
Thus far, several works have been dedicated to crystal tensor property prediction. One prominent category of approaches involves ab initio physical simulation techniques, such as density functional theory (DFT) (Petousis et al., 2016). These classical simulation techniques can accurately calculate various material properties, including electronic structures, phonon spectra, and tensor properties. However, they necessitate extensive computational resources due to the complexity of handling a vast number of atoms and electrons in crystal systems (Yan et al., 2024b), hindering their applicability in practice. Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction. These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Xie & Grossman, 2018) or transformers (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Lee et al., 2024; Wang et al., 2024a).
Existing ML methods have primarily focused on predicting single-value properties of crystals. They exhibit promising prediction performance by enforcing the periodic invariance constraint into the model design. However, these methods cannot be directly applied to tensor property prediction as they overlook the inherent anisotropy of the crystal systems, failing to achieve “special” $O(3)$ equivariance (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Wang et al., 2024a). Some recent works attempt to ensure equivariance through specialized designs of network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023). These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group. In this regard, numerous computationally intensive operations, such as tensor products and the merging of irreducible representations, are required. These operations incur large amounts of computation overhead, particularly when processing high-order data.
To address the above challenges, we propose a novel $O(3)$-equivariant framework for fast crystal tensor property prediction dubbed GoeCTP. In particular, instead of enforcing the equivariance into the model design, we craft a simple yet effective “rotation and reflection” (R&R) module based on polar decomposition. Our designed R&R module rotates and reflects the input crystal with different positions into a standardized invariant position in space. This standardized crystal position is subsequently passed into the property prediction network to obtain an invariant tensor value. Meanwhile, the orthogonal matrix obtained from the R&R module is used to achieve equivariant tensor properties prediction. Our approach is plug-and-play as it can be readily integrated with any existing single-value property prediction network for predicting tensor properties without incurring additional computational costs. Compared to the previous state-of-the-art work Yan et al. (2024b), the GoeCTP method achieves higher quality prediction results and runs more than 13$\times$ faster in the elastic benchmarking dataset.
2 Preliminaries and problem statement
2.1 Preliminaries
The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell. An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters (Yan et al., 2022; Wang et al., 2024b; Jiao et al., 2024). General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
Cartesian Coordinate System. A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{T}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell. Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom. The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{T}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell. The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space. A complete crystal is therefore represented as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\bm{a}_{i}},\hat{\bm{x}}_{i})|\hat{\bm{x}}_{i}=\bm{x}_{i}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3},\hat{\bm{a}_{i}}=\bm{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$. In this representation, the integers $k_{i}$ and $l_{i}$ denote all possible atomic positions in the periodic lattice.
Fractional Coordinate System. Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions. With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{T}\in[0,1)^{3}$. The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$. Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{T}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA (Momma & Izumi, 2011), Yan et al. (2024b), and VELAS (Ran et al., 2023).
2.2 Problem statement
Crystal Tensor Prediction. The crystal tensor properties prediction is a classic regression task. Its goal is to estimate the high-order tensor property denoted as $\mathcal{Y}_{label}$ from the raw crystal data represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$ while minimizing the expected error between the actual property $\mathcal{Y}_{label}$ and the estimated property value $\mathcal{Y}_{pred}$. This problem can be formally formulated as follows:
$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}_{label}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}^{(n)}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset. In what follows, we will omit superscript $n$ for simplicity. As did in the literature (Mao et al., 2024; Wen et al., 2024; Yan et al., 2024b), our objective is to estimate high-order tensor properties, including dielectric tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$) and elastic tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
$\mathbf{O(3)}$-Equivariance. The $O(3)$ group, also known as the orthogonal group in 3D, consists of all rotations and reflections in 3D space, i.e., $3\times 3$ orthogonal matrices. In predicting the crystal tensor properties, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies (Hoogeboom et al., 2022; Song et al., 2024). Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq. 1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
$f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{T},$ (2)
where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix (Yan et al., 2024b). For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig. 1.
Our Objective. Our objective is to devise a new framework that can capture the $O(3)$-equivariance for accurately predicting the high-order crystal tensor properties.
Figure 2: The Illustration of GoeCTP. To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary direction, to the standardized crystal position. Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph. Subsequently, (4) the Prediction module leverages these embedded features to predict the invariant tensor properties corresponding to the standardized crystal position. Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the output of the equivariant tensor properties.
3 Methodology
3.1 The Rationale of Our Framework
To enforce $O(3)$ equivariance, as described in Eq. 2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead. For example, existing work (Yan et al., 2024b; a) gather the rotational information from neighboring nodes to the central node $i$ through the following computation formulated:
$\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})),$ (3)
where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ (Geiger & Smidt, 2022), and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics. When transmitting rotational information for higher-order tensors like the elastic tensor, the rotation order $\lambda$ must be set to 4 (Yan et al., 2024b). In comparison, for transmitting lower-order tensor rotational information (where $\lambda$ is set to 1 (Yan et al., 2024a)), the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ increases fourfold, thus, the corresponding computational cost increases. If we can transmit the rotational information for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, this would significantly reduce the computational costs.
To this end, we propose the GoeCTP method, which achieves $O(3)$ equivariance without requiring specialized network architecture design. For instance, it enables the transmission of rotational information for higher-order tensors using $\lambda=1$ rather than $\lambda=4$, thereby simplifying the model complexity while maintaining equivariance. GoeCTP leverages polar decomposition, a mathematical technique with significant geometric properties. Polar decomposition decomposes a matrix into two components: an orthogonal matrix, representing a rotation or reflection, and a positive semi-definite matrix (Higham, 1986). The orthogonal matrix captures an $O(3)$-group transformation, while the positive semi-definite matrix represents a stretching or scaling. When applied to crystal structures, polar decomposition effectively separates the spatial directional information of the crystal via the orthogonal matrix, allowing us to directly transfer equivariance without the need for intricate model design. Furthermore, polar decomposition is only performed during the data preprocessing stage and does not occur during network training, thereby introducing virtually no additional computational burden.
3.2 Our Proposed Framework: GoeCTP
In what follows, we will first introduce the core rotation and reflection (R&R) module of the proposed GoeCTP, focusing on how to obtain a standardized crystal position for a crystal with arbitrary spatial direction using polar decomposition. Then, we will describe how the input crystal data is processed and introduce the property prediction network of GoeCTP. Finally, we will explain how proposed GoeCTP achieves equivariant property predictions. An overview of the GoeCTP framework is illustrated in Fig. 2.
R&R Module. The primary function of the R&R Module is to rotate and reflect the crystal to a standardized crystal position in space. In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for the R&R module to perform rotation and reflection operations on the crystal data. Therefore, R&R module adopts the fractional coordinate system to represent the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. This choice simplifies the implementation of the R&R module. Specifically, the application of $O(3)$-group transformations is achieved by applying polar decomposition as follows.
Proposition 3.1
(Polar Decomposition (Hall & Hall, 2013; Higham, 1986; Jiao et al., 2024).) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\exp(\mathbf{S})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{S}\in\mathbb{R}^{3\times 3}$ is a symmetric matrix and $\exp(\mathbf{S})=\sum_{n=0}^{\infty}\frac{\mathbf{S}^{n}}{n!}$ defines the exponential mapping of $\mathbf{S}$.
Building on the above proposition, it is evident that for a crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the invertible lattice matrix $\mathbf{L}$ can be uniquely represented by a symmetric matrix $\mathbf{S}$. By defining $\mathbf{H}=\exp(\mathbf{S})$, the lattice matrix $\mathbf{L}$ can also be uniquely represented by a symmetric matrix denoted as $\mathbf{H}$, i.e. $\mathbf{L}=\mathbf{Q}\mathbf{H}$. Furthermore, $\mathbf{H}$ remains invariant with the $O(3)$ group transformation.
Proof 3.1.1
Given the lattice matrices $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$, where $\mathbf{Q^{\prime}}$ is an arbitrary $O(3)$ group transformation, applying polar decomposition on both $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$ yields the following expressions:
$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (4)
$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (5)
By combining Eq. 4 and 5, we can have the following equation:
$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (6)
Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{T}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{T}\mathbf{Q^{\prime}}^{T}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix. According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$. If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition. Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation.
In essence, regardless of the $O(3)$ group transformation applied to the crystal in space, the corresponding $\mathbf{H}$ associated with $\mathbf{L}$ remains invariant. Consequently, $\mathbf{H}$ serves as a fixed standardized crystal position for the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, allowing the crystal with different position to be consistently rotated to a same position through polar decomposition. The concept of the standardized crystal position can be formally stated as the following proposition :
Proposition 3.2
(Standardized Crystal Position.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$, the standardized crystal position $\mathbf{H}$ is obtained. When an arbitrary $O(3)$ group transformation $\mathbf{Q^{\prime}}$ is applied to a crystal structure, $\mathbf{H}$ remains unchanged.
Therefore, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ formulated as:
$f_{p}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H}),$ (7)
where $f_{p}(\mathbf{M})$ represents the application of polar decomposition to the lattice matrix $\mathbf{L}$ of $\mathbf{M}$.
As shown in Fig. 2, the crystal $(\mathbf{A},\mathbf{F},\mathbf{H})$ is passed to the subsequent crystal graph construction module for further processing. The orthogonal matrix $\mathbf{Q}$ obtained during this decomposition is passed to the reverse R&R module, ensuring the equivariant transformation of the output tensor properties. Our proposed R&R module based on polar decomposition allows the input crystal data to be transformed into a standardized spatial position that is invariant under $O(3)$ space group transformations. With this particular module, the equivariance can be captured, meaning that the subsequent components of GeoCTP are no longer required to account for equivariance.
Crystal Graph Construction. To enable networks to handle such infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is typically necessary to employ graph construction methods that represent the interactions between infinite crystal structures and atoms using finite graph data. Here, we use the crystal graph construction from Yan et al. (2024a; b) to describe the structure and relationships within crystals. Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{P},\mathbf{H})$ as introduced in Sec. 2. We assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$. $\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$. When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e. $d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$. Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$. Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al. (2024a); Yan et al. (2024a).
Node and Edge Feature Embedding. Building on previous work (Xie & Grossman, 2018; Yan et al., 2024b; a), node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$. Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$. The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding (Lin et al., 2023). Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
Prediction Module. Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al. (2024a); Taniai et al. (2024); Yan et al. (2022); Lee et al. (2024); Wang et al. (2024a), among others. For better performance, we select eComformer (Yan et al., 2024a), which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module. A detailed explanation of eComFormer can be found in Appendix A.1. Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows:
$\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}.$ (8)
Reverse R&R Module. The primary function of the Reverse R&R module is to generate the equivariant tensor property predictions based on the crystal global features from the Prediction module. First, the Reverse R&R module transforms $\bm{G}^{\mathrm{final}}$ into a tensor output, as follows:
$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}}),$ (9)
$\bm{\varepsilon}^{\mathrm{final}}=f_{rp}(\bm{\varepsilon},\mathbf{Q}),$ (10)
where $f_{MLP}(\cdot)$ represents a multilayer perceptron (MLP) and the operation reshaping dimension. Next, $f_{rp}(\cdot)$ utilizes the orthogonal matrix $\mathbf{Q}$ obtained from the R&R Module to convert the tensor output $\bm{\varepsilon}$ into its final equivariant form denoted as $\bm{\varepsilon}^{\mathrm{final}}$.
This conversion $f_{rp}(\cdot)$ for predicting the dielectric tensor can be expressed by:
$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{T}.$ (11)
For predicting the higher-order elastic tensor, the conversion process becomes more complex. Specifically, the conversion is expressed as follows:
$C_{ijkl}^{\mathrm{final}}=\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{p=1}^{3}\sum_{q=1}^{3}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq},$ (12)
where the subscripts denote the specific positions of the elements within the tensor data.
4 Related work
GNN-Based Methods. CGCNN is a pioneering GNN model specifically designed for handling crystal structures (Xie & Grossman, 2018). This model proposed to represent crystal structures as multi-edge crystal graphs. It was applied to predict various single-value properties such as formation energy and band gap. Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Das et al., 2023; Lin et al., 2023). These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors. Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction. In contrast, recent studies attempted to ensure equivariance through specialized network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023). These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties. Within these network architectures, many operations are required, such as tensor products and combining irreducible representations. These processes significantly increase computational costs, especially when handling higher-order data.
Transformer-Based Methods. Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties. Matformer (Yan et al., 2022), one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells. This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures. Subsequently, more advanced Transformer-based approaches were proposed. Some methods, such as ComFormer (Yan et al., 2024a), CrystalFormer (Wang et al., 2024a), and CrystalFormer (Wang et al., 2024a), typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties. Furthermore, DOSTransformer (Lee et al., 2024) is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance. Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
5 Experiments
5.1 Experimental setup
Datasets. In this work, we evaluate the performance of GoeCTP on two key tensor property prediction tasks: the second-order dielectric tensor and the fourth-order elastic tensor, respectively. The dataset for dielectric tensor is derived from the data processed by Yan et al. (2024b), sourced from the JARVIS-DFT database. Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/. The statistical details of the datasets are presented in Table 1. Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix. Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.2 for details) and then reconstruct the entire $3\times 3$ symmetric matrix. In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor (Itin & Hehl, 2013; Wen et al., 2024). Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.2 for details).
Dataset Sample size Fnorm Mean Fnorm STD Unit
Dielectric 4713 14.7 18.2 Unitless
Elastic 25110 306.4 238.4 GPa
Table 1: Dataset statistics.
Baseline Methods. we selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET (Chen et al., 2019; Morita et al., 2020), EGTNN (Zhong et al., 2023), and GMTNet (Yan et al., 2024b), as baseline methods.
[…]
Experimental Settings. The experiments were performed using an NVIDIA GeForce RTX 3090 GPU. For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al. (2024b). For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. During model training, Huber loss, AdamW (Loshchilov & Hutter, 2018), a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
Since we observed some randomness in the results during the training process of the repeated dielectric dataset experiments, we conducted 5 repeated experiments for each method and used the average of the 5 experiment metrics as the final experimental results. The results of the 5 repeated experiments can be found in the Appendix 11.
5.2 Experimental Results
Predicting Dielectric Tensors. The performance of various models in predicting the dielectric tensor is summarized in Table 2. While GoeCTP and GMTNet show close results in terms of Fnorm and EwT 25%, GoeCTP demonstrates higher values for EwT 10% and EwT 5%, with an improvement of 5% in EwT 5% compared to GMTNet. This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet. Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024). Detailed results of these combinations are provided in Appendix 11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5%
Table 2: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the dielectric dataset.
Predicting Elastic Tensors. Experimental results on the elastic tensor dataset are presented in Table 3. GoeCTP demonstrated outstanding performance in predicting higher-order, complex tensors, outperforming all baseline methods across every evaluation metric. Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science. Further details on the results of GoeCTP’s combination with other models can be found in Appendix 11.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2%
Table 3: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the elastic dataset.
Verifying the $O(3)$ Equivariance. To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties. Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set). All crystals in the original test set were adjusted to the standardized crystal position, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set. The method for generating the corresponding orthogonal matrices is from Heiberger (1978). We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
The results on the dielectric tensor dataset are shown in Table 4. GoeCTP performed equally well on the augmented test set as on the original test set, indicating that it maintains strong $O(3)$ equivariance for tensor properties. In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45%, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties. The results on the elastic tensor dataset are shown in Table 5. Similar to the dielectric tensor dataset, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm dropping by nearly 29%. In contrast, GoeCTP’s performance remained unchanged, further demonstrating the effectiveness of the GoeCTP method. Additionally, runtime comparisons between GoeCTP and eComFormer on the test set showed that GoeCTP introduced no significant increase in runtime. This indicates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality without compromising efficiency.
eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
EwT 25% $\uparrow$ 83.2% 69.7% 83.2% 83.2%
EwT 10% $\uparrow$ 56.8% 42.7% 56.8% 56.8%
EwT 5% $\uparrow$ 35.5% 22.5% 35.5% 35.5%
Total Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric dataset.
eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
EwT 25% $\uparrow$ 42.5% 25.9% 42.5% 42.5%
EwT 10% $\uparrow$ 15.3% 2.2% 15.3% 15.3%
EwT 5% $\uparrow$ 7.2% 0.2% 7.2% 7.2%
Total Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
Table 5: Ablation study for verifying the $O(3)$ equivariance with elastic dataset.
Efficiency. The results presented in Table 6 and Table 7 illustrate the running times for GoeCTP and baseline methods. On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet. On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process. To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply. In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a multilayer perceptron (MLP) at the network’s output, ensuring both efficiency and scalability across different tensor orders.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 663 1325 1611 616
Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048
Table 6: Efficiency comparison on the dielectric dataset.
MEGNET ETGNN GMTNet GoeCTP (Ours)
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422
Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189
Table 7: Efficiency comparison on the elastic dataset.
6 Conclusion, limitations, and future works
In this work, we propose a novel $O(3)$-equivariant framework GoeCTP for fast crystal tensor prediction, which is a plug-and-play framework that can be integrated with any existing single-value property prediction network, enabling them to predict tensor properties with almost no additional computational cost. Based on predicting invariant tensor properties from standardized crystal positions and using an external module to ensure tensor equivariance, this approach has achieved state-of-the-art performance and the highest efficiency in different widely used crystal tensor property benchmark tests. The limitations of our current GoeCTP include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module and, in theory, does not introduce additional performance enhancements. (2) Currently, GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types. In future work, we intend to explore the following directions: for (1), we aim to improve performance by integrating prior knowledge related to the independent components of tensor properties across different crystal systems in Appendix A.2; for (2), we plan to extend the framework to other domains, such as 3D point cloud research, broadening its applicability.
Appendix A Appendix
A.1 Details of eComFormer
The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks (Yan et al., 2024a). It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph. The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. (2024a).
[…]
$\bm{\alpha}_{ij}=\frac{\bm{q}_{ij}\circ\bm{\xi}_{K}(\bm{k}_{ij})}{\sqrt{d_{\bm{q}_{ij}}}},\bm{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\bm{\alpha}_{ij}))\circ\bm{\xi}_{V}(\bm{\upsilon}_{ij}),$ (13)
where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively, and the operators $\circ$ denote the Hadamard product, $\bm{\xi}_{msg}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$. Then, node feature $\bm{f}_{i}$ is updated as follows,
$\bm{m}s\bm{g}_{i}=\sum_{j\in\mathcal{N}_{i}}\bm{m}s\bm{g}_{ij},\bm{f}_{i}^{\mathrm{new}}=\bm{\xi}_{msg}(\bm{f}_{i}+\mathrm{BN}(\bm{m}s\bm{g}_{i})),$ (14)
where $\bm{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers (Geiger & Smidt, 2022). It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features. Gathering rotational information from neighboring nodes to the central node $i$, the first TP layer is shown as
$\bm{f}_{i,0}^{l}=\bm{f}_{i}^{l^{\prime}}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{0}(\hat{\mathbf{e_{ji}}})),\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})),\lambda\in\{1,2\},$ (15)
where $\bm{f}_{i}^{l^{\prime}}$ is linearly transformed from $\bm{f}_{i}^{l}$, $|\mathcal{N}_{i}|$ and $\mathbf{TP}_{\lambda}$ represent the number of neighbors of node $i$ and TP layers with rotation order $\lambda$ respectively. Then, to represent the invariant node features, the second TP layer is written as
$\bm{f}_{i}^{l*}=\frac{1}{|\mathcal{N}_{i}|}(\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,0}^{l},\mathbf{Y}_{0}(\hat{\mathbf{e_{ji}}}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,1}^{l},\mathbf{Y}_{1}(\hat{\mathbf{e_{ji}}}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,2}^{l},\mathbf{Y}_{2}(\hat{\mathbf{e_{ji}}}))),$ (16)
stacking the two tensor product layers together using both linear and nonlinear transformations, the output $\bm{f}_{i,updated}^{l}$ is combined as
$\bm{f}_{i,updated}^{l}=\bm{\sigma}_{\mathrm{equi}}(\mathrm{BN}(\bm{f}_{i}^{l*}))+\mathrm{LN}_{\mathrm{equi}}(\bm{f}_{i}^{l}),$ (17)
with $\bm{\sigma}_{\mathrm{equi}}$ denoting a nonlinear transformation made up of two softplus layers with a linear layer positioned between them. The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer are shown in Fig. 3.
A.2 Tensor properties symmetry
Tensor properties, such as dielectric tensors and elastic tensors, describe the material’s response to external physical fields (such as electric fields or mechanical stress). In materials with symmetry, this response must adhere to the symmetry requirements of the material. As demonstrated by Yan et al. (2024b), when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$,the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$. Therefore, the corresponding tensor properties also remain unchanged, i.e., $\bm{\varepsilon}=\mathbf{R}\bm{\varepsilon}\mathbf{R}^{T}$. Thus, crystal symmetry imposes strict constraints on the components of the tensor, leading to the simplification or elimination of many components, reducing the number of independent components.
Crystal system Number of independent elements Dielectric tensor
Cubic 1 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
Tetragonal & Hexagonal & Trigonal 2 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Orthorhombic 3 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{22}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Monoclinic 4 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&\bm{\varepsilon}_{13}\\ 0&\bm{\varepsilon}_{22}&0\\ \bm{\varepsilon}_{13}&0&\bm{\varepsilon}_{33}\end{pmatrix}$
Triclinic 6 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&\bm{\varepsilon}_{12}&\bm{\varepsilon}_{13}\\ \bm{\varepsilon}_{12}&\bm{\varepsilon}_{22}&\bm{\varepsilon}_{23}\\ \bm{\varepsilon}_{13}&\bm{\varepsilon}_{23}&\bm{\varepsilon}_{33}\end{pmatrix}$
Table 8: Number of independent components in the dielectric tensor for different crystal systems.
Independent components in the dielectric tensor. The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry. The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 8 (Mao et al., 2024).
Voigt notation for elastic tensor. Similar to the dielectric tensor, the elastic tensor also has independent elements for various types of systems due to the crystal symmetry. The elastic tensor has a minimum of 3 and a maximum of 21 independent elements for various types of systems. Voigt notation is a compact way to represent these independent components of tensor properties (Itin & Hehl, 2013). According to the rules $11\to 1\mathrm{~{};~{}}22\to 2\mathrm{~{};~{}}33\to 3\mathrm{~{};~{}}23,~{}32\to 4\mathrm{~{};~{}}31,~{}13\to 5\mathrm{~{};~{}}12,~{}21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix (Wen et al., 2024; Ran et al., 2023):
$C=\begin{pmatrix}C_{1111}&C_{1122}&C_{1133}&C_{1123}&C_{1131}&C_{1112}\\ C_{1122}&C_{2222}&C_{2233}&C_{2223}&C_{2231}&C_{2212}\\ C_{1133}&C_{2233}&C_{3333}&C_{3323}&C_{3331}&C_{3312}\\ C_{1123}&C_{2223}&C_{3323}&C_{2323}&C_{2331}&C_{2312}\\ C_{1131}&C_{2231}&C_{3331}&C_{2331}&C_{3131}&C_{3112}\\ C_{1112}&C_{2212}&C_{3312}&C_{2312}&C_{3112}&C_{1212}\end{pmatrix}\to\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}$
The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to Wen et al. (2024); Ran et al. (2023)).
Crystal system Number of independent elements Elastic tensor
Cubic 3 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{12}&0&0&0\\ C_{12}&C_{11}&C_{12}&0&0&0\\ C_{12}&C_{12}&C_{11}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{44}&0\\ 0&0&0&0&0&C_{44}\end{pmatrix}\end{gathered}$
Tetragonal 6 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&0&0&0\\ C_{12}&C_{11}&C_{13}&0&0&0\\ C_{13}&C_{13}&C_{33}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{44}&0\\ 0&0&0&0&0&C_{66}\end{pmatrix}\end{gathered}$
Orthorhombic 9 $C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&0&0&0\\ C_{12}&C_{22}&C_{23}&0&0&0\\ C_{13}&C_{23}&C_{33}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{55}&0\\ 0&0&0&0&0&C_{66}\end{pmatrix}$
Triclinic 21 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}\end{gathered}$
Table 9: Number of independent components in the elastic tensor for different crystal systems.
A.3 Experimental details and additional results
Hyperparameter settings of GoeCTP. When constructing the crystal graph, we used the 16th nearest atom to determine the cutoff radius. For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector. In the dielectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector. For the prediction module, eComFormer, in the dielectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer. For both dielectric and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64. In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output. For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
1st 2nd 3rd 4th 5th mean
MEGNET
Fnorm $\downarrow$ 3.62 3.70 3.66 3.74 3.81 3.71
EwT 25% $\uparrow$ 76.0% 77.1% 75.6% 73.7% 76.4% 75.8 %
EwT 10% $\uparrow$ 41.2% 39.3% 38.2% 39.7% 36.3% 38.9%
EwT 5% $\uparrow$ 17.6% 16.8% 18.5% 19.1% 17.8% 18.0%
ETGNN
Fnorm $\downarrow$ 3.42 3.48 3.33 3.35 3.41 3.40
EwT 25% $\uparrow$ 82.1% 81.7% 83.4% 82.6% 83.2% 82.6 %
EwT 10% $\uparrow$ 46.9% 47.8% 50.7% 49.4% 50.5% 49.1%
EwT 5% $\uparrow$ 22.3% 26.5% 27.1% 25.9% 24.6% 25.3%
GMTNet
Fnorm $\downarrow$ 2.98 3.64 3.23 3.16 3.39 3.28
EwT 25% $\uparrow$ 83.4% 81.3% 83.8% 84.5% 83.7% 83.3 %
EwT 10% $\uparrow$ 54.9% 56.2% 55.8% 56.2% 57.1% 56.0%
EwT 5% $\uparrow$ 31.2% 30.5% 28.8% 29.3% 32.7% 30.5%
GoeCTP (Ours)
Fnorm $\downarrow$ 3.33 3.29 3.36 3.25 2.92 3.23
EwT 25% $\uparrow$ 82.5% 83.4% 82.1% 83.9% 84.2% 83.2 %
EwT 10% $\uparrow$ 57.5% 53.9% 56.4% 56.7% 59.7% 56.8%
EwT 5% $\uparrow$ 36.9% 31.6% 37.1% 34.2% 37.8% 35.5%
Table 10: Comparison of 5 repeated experiments on dielectric dataset.
Hyperparameter settings of GMTNet, ETGNN, and MEGNET. Following Yan et al. (2024b), we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks. The same polynomial learning rate decay scheduler is used in all experiments.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23 3.40 3.53
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2% 81.7% 80.1%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8% 53.8% 52.9%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5% 32.3% 30.6%
Total Time (s) $\downarrow$ 663 1325 1611 616 535 645
Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048 0.042 0.202
Table 11: Additional comparison of performance metrics on dielectric dataset.
Additional results. Given the relatively small size of the dielectric tensor dataset, along with other influencing factors, we observed some randomness in the results during repeated training runs, introducing noise into the outcome comparisons. To mitigate this effect and ensure the reliability of the results, we conducted five independent trials for each method, using the average of the performance metrics across these five trials as the final reported value. The detailed results from these repeated experiments are presented in Table 10.
We assessed the performance of GoeCTP when combined with iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024) (denoted as GoeCTP (iCom.) for the combination) on both the dielectric tensor and elastic tensor datasets. As presented in Table 11, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). In contrast, as shown in Table 12, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11 102.80 107.44
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5% 46.7% 43.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3% 18.6% 15.8%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2% 8.2% 7.9%
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422 4035 7891
Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189 0.315 0.616
Table 12: Additional comparison of performance metrics on elastic dataset.

### v3

@@
-Fast Crystal Tensor Property Prediction: A General O(3)-Equivariant Framework Based on Polar Decomposition
+Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction
 Abstract
@@
 Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance.
-This equivariance constraint often introduces tremendous computational costs, necessitating specialized designs for effective and efficient predictions.
-To address this limitation, we propose a general $O(3)$-equivariant framework for fast crystal tensor property prediction, called GoeCTP.
-Our framework is efficient as it does not need to impose equivalence constraints onto the network architecture.
-Instead, GoeCTP captures the tensor equivariance with a simple external rotation and reflection (R&R) module based on polar decomposition.
-The crafted external R&R module can rotate and reflect the crystal into an invariant standardized crystal position in space without introducing extra computational cost.
-We show that GoeCTP is general as it is a plug-and-play module that can be smoothly integrated with any existing single-value property prediction framework for predicting tensor properties.
-Experimental results indicate that GoeCTP achieves higher prediction performance and runs 13$\times$ faster compared to existing state-of-the-art methods in elastic benchmarking datasets, underscoring its effectiveness and efficiency.
+Such equivariance constraint often requires specialized architecture designs to achieve effective predictions, inevitably introducing tremendous computational costs.
+Canonicalization, a classical technique for geometry, has recently been explored for efficient learning with symmetry.
+In this work, we revisit the problem of crystal tensor property prediction through the lens of canonicalization.
+Specifically, we demonstrate how polar decomposition, a simple yet efficient algebraic method, can serve as a form of canonicalization and be leveraged to ensure equivariant tensor property prediction.
+Building upon this insight, we propose a general $O(3)$-equivariant framework for fast and accurate crystal tensor property prediction, referred to as GoeCTP.
+By utilizing canonicalization, GoeCTP achieves high efficiency without requiring the explicit incorporation of equivariance constraints into the network architecture.
+Experimental results indicate that GoeCTP achieves the best prediction accuracy and runs up to 13$\times$ faster compared to existing state-of-the-art methods in benchmarking datasets, underscoring its effectiveness and efficiency.
 1 Introduction
-The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering (Yan et al., 2024b).
-Compared to single-value property prediction, predicting tensor properties of crystalline materials is substantially more complex.
-This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress (Nye, 1985; Resta, 1994; Yan et al., 2024b).
-Consequently, tensor property prediction modeling necessitates preserving consistency with the crystal’s spatial position, exhibiting “special” $O(3)$ equivariance (Yan et al., 2024b).
-Furthermore, high-order tensor property prediction is computationally intensive due to its high dimensionality and large data volume.
-Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
+The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering [51].
+These tensor properties span various orders, such as dielectric tensor with two orders, piezoelectric tensor with three orders, and elastic tensor with four orders.
+Accurate prediction of these tensor properties is critical for novel materials discovery and design with targeted characteristics.
 Thus far, several works have been dedicated to crystal tensor property prediction.
-One prominent category of approaches involves ab initio physical simulation techniques, such as density functional theory (DFT) (Petousis et al., 2016).
-These classical simulation techniques can accurately calculate various material properties, including electronic structures, phonon spectra, and tensor properties.
-However, they necessitate extensive computational resources due to the complexity of handling a vast number of atoms and electrons in crystal systems (Yan et al., 2024b), hindering their applicability in practice.
+One prominent category involves ab initio physical simulation techniques, such as density functional theory (DFT) [35].
+These classical simulation techniques can provide acceptable error margin for predicting various material properties.
+However, they necessitate extensive computational resources due to the complexity of handling large crystals with a vast number of atoms and electrons [51], hindering their applicability in practice.
 Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction.
-These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Xie & Grossman, 2018) or transformers (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Lee et al., 2024; Wang et al., 2024a).
-Existing ML methods have primarily focused on predicting single-value properties of crystals.
-They exhibit promising prediction performance by enforcing the periodic invariance constraint into the model design.
-However, these methods cannot be directly applied to tensor property prediction as they overlook the inherent anisotropy of the crystal systems, failing to achieve “special” $O(3)$ equivariance (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Wang et al., 2024a).
-Some recent works attempt to ensure equivariance through specialized designs of network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023).
+These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) [1, 26, 2, 47] or transformers [50, 42, 49, 20, 44].
+Most existing methods are designed for scalar property prediction, focusing on achieving $SO(3)$ invariance of the crystal structures.
+However, these scalar-property methods could not be adapted to predict tensor properties of crystalline materials, which exhibits significantly higher complexity.
+This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress [33, 38, 51]; their modeling necessitates preserving consistency with the crystal’s spatial orientation, exhibiting a unique tensor $O(3)$ equivariance [51].
+Therefore, a few recent studies attempt to ensure equivariance through specialized designs of network architectures [28, 25, 11, 46, 34, 51, 53].
 These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group.
@@
 These operations incur large amounts of computation overhead, particularly when processing high-order data.
-To address the above challenges, we propose a novel $O(3)$-equivariant framework for fast crystal tensor property prediction dubbed GoeCTP.
-In particular, instead of enforcing the equivariance into the model design, we craft a simple yet effective “rotation and reflection” (R&R) module based on polar decomposition.
-Our designed R&R module rotates and reflects the input crystal with different positions into a standardized invariant position in space.
-This standardized crystal position is subsequently passed into the property prediction network to obtain an invariant tensor value.
-Meanwhile, the orthogonal matrix obtained from the R&R module is used to achieve equivariant tensor properties prediction.
-Our approach is plug-and-play as it can be readily integrated with any existing single-value property prediction network for predicting tensor properties without incurring additional computational costs.
+Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
+One approach that has show promising results in efficient learning with symmetry across various fields, including point clouds [22], n-body simulation [17], and antibodies generation [29], is the use of “canonicalization".
+Specifically, canonicalization maps a geometric data to an invariant representation [27, 6], referred to as the canonical form, and subsequently reconstructs an equivariant output from the canonical form without imposing any architectural constraints on the network.
+However, to date, it has not been explored for crystal tensor property prediction.
+In this work, we revisit the task of crystal tensor property prediction through the lens of canonicalization and introduce a novel canonicalization strategy tailored for this particular setting.
+Specifically, rather than embedding equivariance directly into model architecture, we propose a simple yet effective canonicalization module, termed R&R.
+In particular, our canonicalization is instantiated based on polar decomposition, a continuous mapping technique [6] that can provide enhanced robustness.
+During prediction, the R&R module applies rotations and reflections to transform the input crystal structure into its canonical form.
+The canonical form is then fed into a property prediction network to obtain the corresponding canonical tensor representation.
+Simultaneously, the orthogonal matrix derived from the R&R module is utilized to recover the equivariant output via the tensor transformation rule, enabling equivariant tensor property prediction without incurring additional computational overhead.
+We conducted experiments on dielectric, piezoelectric, and elastic tensor datasets, respectively, showcasing that the proposed method can achieve $O(3)$-equivariant tensor prediction while maintaining high efficiency.
 Compared to the previous state-of-the-art work Yan et al.
-(2024b), the GoeCTP method achieves higher quality prediction results and runs more than 13$\times$ faster in the elastic benchmarking dataset.
-2 Preliminaries and problem statement
+[51], the GoeCTP method achieves higher prediction accuracy and runs by up to 13$\times$ faster in benchmarking datasets.
+2 Background
 2.1 Preliminaries
 The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell.
-An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters (Yan et al., 2022; Wang et al., 2024b; Jiao et al., 2024).
+An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters [49, 45, 16].
 General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
 Cartesian Coordinate System.
-A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{T}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell.
+A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{\top}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell.
 Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom.
-The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{T}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell.
+The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell.
 The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space.
@@
 Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions.
-With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{T}\in[0,1)^{3}$.
+With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{\top}\in[0,1)^{3}$.
 The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$.
-Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{T}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
-Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction.
-Some of the visualizations in the figure are generated using the VESTA (Momma & Izumi, 2011), Yan et al.
-(2024b), and VELAS (Ran et al., 2023).
+Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
+Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA [30], Yan et al. [51], and VELAS [37].
 2.2 Problem statement
@@
 The crystal tensor properties prediction is a classic regression task.
-Its goal is to estimate the high-order tensor property denoted as $\mathcal{Y}_{label}$ from the raw crystal data represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$ while minimizing the expected error between the actual property $\mathcal{Y}_{label}$ and the estimated property value $\mathcal{Y}_{pred}$.
-This problem can be formally formulated as follows:
-$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}_{label}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}^{(n)}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
+Its goal is to estimate the high-order tensor property denoted as $\mathcal{Y}$ from the raw crystal data represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$.
+Given that the crystal data $\mathbf{M}$ resides within the input space $\mathcal{V}$, and the tensor property $\mathcal{Y}$ belongs to the separate output space $\mathcal{W}$, the objective of crystal tensor prediction is to find a function ${f_{\theta}}:\mathcal{V}\to\mathcal{W}$ that accurately maps input crystal data to the desired tensor property.
+This is achieved by minimizing the discrepancy between the true property $\mathcal{Y}$ and the predicted property value $\mathcal{Y}_{pred}$.
+Therefore, the crystal tensor property prediction task can be mathematically formulated as the following optimization problem:
+$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
 where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset.
 In what follows, we will omit superscript $n$ for simplicity.
-As did in the literature (Mao et al., 2024; Wen et al., 2024; Yan et al., 2024b), our objective is to estimate high-order tensor properties, including dielectric tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$) and elastic tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
-$\mathbf{O(3)}$-Equivariance.
-The $O(3)$ group, also known as the orthogonal group in 3D, consists of all rotations and reflections in 3D space, i.e., $3\times 3$ orthogonal matrices.
-In predicting the crystal tensor properties, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies (Hoogeboom et al., 2022; Song et al., 2024).
+As did in the literature [28, 46, 51], our objective is to estimate high-order tensor properties, including 2-order dielectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$), 3-order piezoelectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$), and 4-order elastic tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
+$\mathbf{O(3)}$ group.
+The $O(3)$ group is an orthogonal group, consisting of rotations and reflections.
+The elements $g\in O(3)$ act on vectors or tensors in $\mathcal{V}$, $\mathcal{W}$ through their respective group representations, $\rho_{\mathcal{V}}:O(3)\to GL(\mathcal{V})$ and $\rho_{\mathcal{W}}:O(3)\to GL(\mathcal{W})$, where $GL(\mathcal{V})$ and $GL(\mathcal{W})$ are the space of invertible linear maps $\mathcal{V}\to\mathcal{V}$ and $\mathcal{W}\to\mathcal{W}$, respectively.
+Specifically, the action of an element $g\in O(3)$ on crystal data $\mathbf{M}$ is defined as: $g\cdot\mathbf{M}=\rho_{\mathcal{V}}(g)\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})$.
+For dielectric tensor property $\mathcal{Y}$, the action of $g$ is given by: $g\cdot\mathcal{Y}=\rho_{\mathcal{W}}(g)\mathcal{Y}=\mathbf{Q}\mathcal{Y}\mathbf{Q}^{\top}$, where $\mathbf{Q}$ is a $3\times 3$ orthogonal matrix (for additional details regarding the transformation of tensor properties, please refer to Appendix A.8).
+$\mathbf{O(3)}$-Equivariance for Crystal Tensor Prediction.
+In the crystal tensor properties prediction task, due to the difference between $\rho_{\mathcal{V}}(g)$ and $\rho_{\mathcal{W}}(g)$, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies [13, 48, 52, 40].
 Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq.
 1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
-$f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{T},$ (2)
-where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix (Yan et al., 2024b). For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig. 1.
-Our Objective. Our objective is to devise a new framework that can capture the $O(3)$-equivariance for accurately predicting the high-order crystal tensor properties.
+$f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top},$ (2)
+where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix [51].
+For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig.
+1 (for more equivariance details, see Appendix A.8).
 Figure 2: The Illustration of GoeCTP.
-To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary direction, to the standardized crystal position.
+To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary orientation, to the canonical form of this crystal.
 Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph.
-Subsequently, (4) the Prediction module leverages these embedded features to predict the invariant tensor properties corresponding to the standardized crystal position.
-Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the output of the equivariant tensor properties.
+Subsequently, (4) the Prediction module leverages these embedded features to predict the canonical form of tensor properties corresponding to the canonical form of input crystal.
+Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the equivariance of the output tensor properties.
 3 Methodology
-3.1 The Rationale of Our Framework
+In this section, we will first provide the motivation driving our new framework.
+We then reformulate the $O(3)$ equivariant tensor prediction task from the perspective of canonicalization.
+Building upon this foundational perspective, we demonstrate how polar decomposition can be employed as a canonicalization strategy and effectively applied to solve tensor prediction tasks with both effectiveness and efficiency.
+3.1 The Motivation of Our Framework
 To enforce $O(3)$ equivariance, as described in Eq.
 2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead.
-For example, existing work (Yan et al., 2024b; a) gather the rotational information from neighboring nodes to the central node $i$ through the following computation formulated:
+For example, existing work gather the rotational information from the neighboring nodes to the central node $i$ through the following computation formulated [51]:
 $\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),$ (3)
-where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ (Geiger & Smidt, 2022), and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics.
-When transmitting rotational information for higher-order tensors like the elastic tensor, the rotation order $\lambda$ must be set to 4 (Yan et al., 2024b).
-In comparison, for transmitting lower-order tensor rotational information (where $\lambda$ is set to 1 (Yan et al., 2024a)), the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ increases fourfold, thus, the corresponding computational cost increases.
-If we can transmit the rotational information for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, this would significantly reduce the computational costs.
-To this end, we propose the GoeCTP method, which achieves $O(3)$ equivariance without requiring specialized network architecture design.
-For instance, it enables the transmission of rotational information for higher-order tensors using $\lambda=1$ rather than $\lambda=4$, thereby simplifying the model complexity while maintaining equivariance.
-GoeCTP leverages polar decomposition, a mathematical technique with significant geometric properties.
-Polar decomposition decomposes a matrix into two components: an orthogonal matrix, representing a rotation or reflection, and a positive semi-definite matrix (Higham, 1986).
-The orthogonal matrix captures an $O(3)$-group transformation, while the positive semi-definite matrix represents a stretching or scaling.
-When applied to crystal structures, polar decomposition effectively separates the spatial directional information of the crystal via the orthogonal matrix, allowing us to directly transfer equivariance without the need for intricate model design.
-Furthermore, polar decomposition is only performed during the data preprocessing stage and does not occur during network training, thereby introducing virtually no additional computational burden.
+where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ [7], and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics.
+To achieve $O(3)$ equivariance of the model and ensure high prediction accuracy for high-order tensor prediction tasks, such as elastic tensor prediction, the rotation order $\lambda$ is set to 3 [51], i.e., $\mathbf{Y}_{3}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{7}$.
+In contrast, for lower-order tensor prediction tasks, setting $\lambda$ to 1 is sufficient to achieve the model’s $O(3)$ equivariance, i.e., $\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{3}$.
+Compared to lower-order tensor prediction, the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ nearly doubles for high-order tensor predictions, leading to a significant increase in computational complexity.
+Consequently, this model architecture introduces tremendous computational costs for tensor property predictions.
+If we can achieve the model’s $O(3)$ equivariance and maintain the model performance for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, or even without relying on harmonic decomposition, computational costs could be drastically reduced.
+Inspired by the recent exploration of using canonicalization in efficient learning with symmetry, we aim to propose a principled approach based on canonicalization to enable $O(3)$-equivariant tensor predictions for crystal materials.
+3.2 $O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
+To mathematically describe $O(3)$-equivariant tensor prediction from the perspective of canonicalization, we first introduce and extend several fundamental concepts derived from canonicalization techniques [27, 17, 6].
+Definition 3.1.
+(Orbit.) The orbit of a crystal $\mathbf{M}$ is defined as $\text{Orbit}(\mathbf{M})=\{g\cdot\mathbf{M}\mid g\in O(3)\}$.
+In the context of crystal tensor prediction, all configurations of a crystal (i.e.
+a crystal wirh different orientations) in the space fall into the same orbit under the $O(3)$ group transformation.
+Specifically, for any two crystal configurations denoted as $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ within the same orbit, there exists a rotation or reflection $g\in O(3)$ such that $\mathbf{M}_{1}=g\cdot\mathbf{M}_{2}$.
+Likewise, the tensor properties of the crystal also exhibit an orbit, i.e.
+$\text{Orbit}(\mathcal{Y})=\{g\cdot\mathcal{Y}\mid g\in O(3)\}$.
+All configurations of a crystal within an orbit could be transformed into one another through group transformations.
+Consequently, we can select a specific configuration from the orbit to serve as a canonical representative (referred to as the canonical form) that effectively characterizes the entire orbit.
+The formal process is outlined in the following definition.
+Definition 3.2.
+(Orbit Canonicalization.) A function denoted as $C_{\mathbf{M}}:\mathcal{V}\to\mathcal{V}$ is defined as an orbit canonicalization, if $\forall\mathbf{M}_{1}\in\text{Orbit}(\mathbf{M})$ and $\exists\mathbf{M}_{0}\in\text{Orbit}(\mathbf{M}),$ such that $C_{\mathbf{M}}(\mathbf{M}_{1})=\mathbf{M}_{0}$, where $\mathbf{M}_{0}$ is the canonical form of the orbit.
+Orbit canonicalization enables the transformation of a crystal with arbitrary orientations in the space into one with a specific, well-defined orientation.
+Since the output of orbit canonicalization can always be mapped to the invariant canonical form, it follows that $C_{\mathbf{M}}(\cdot)$ itself is $O(3)$-invariant.
+Therefore, for any prediction function ${f_{\theta}}(\cdot)$, it holds that ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ remains $O(3)$-invariant (See the proof in Appendix A.1).
+Although an $O(3)$-invariant prediction function has been established via orbit canonicalization, the complete process of equivariant prediction requires specific group transformations that map all crystal configurations within the orbit to their canonical form.
+To address this requirement, we introduce the concept of rigid registration.
+Rigid registration is a fundamental concept in fields such as computer vision and computational geometry [41].
+It aims to find a transformation that matches the position and orientation of one object with the corresponding position and orientation of another object.
+This concept can be naturally adapted to the crystal, as follows.
+Definition 3.3.
+(Rigid Registration.) A function $R_{\mathbf{M}}:\mathcal{V}\times\mathcal{V}\to O(3)$ is a rigid registration if $\forall\mathbf{M}_{1},\mathbf{M}_{2}\in\text{Orbit}(\mathbf{M}),R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})=g$, such that $\mathbf{M}_{1}=g\cdot\mathbf{M}_{2}$.
+Using Definition 3.3, we can identify a transformation that matches two configurations of a crystal.
+Specifically, this allows us to find specific group transformations that map all crystal configurations within the orbit to their canonical form.
+By leveraging Definition 3.2 and Definition 3.3 together, we derive a view for performing $O(3)$-equivariant tensor prediction through canonicalization, as described below.
+Proposition 3.4.
+($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function formulated as $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$-equivariant for tensor prediction.
+The proof of Proposition 3.4 is provided in Appendix A.6.
+Specifically, in this unified view of $O(3)$ equivariant tensor prediction, we leverage rigid registration, as defined in Definition 3.3, to find specific group transformations that map crystal configurations within the orbit to their canonical form.
+This canonical form is determined through orbit canonicalization, as described in Definition 3.2.
+Subsequently, we can directly apply these transformations to the network’s output, thereby effectively achieving the $O(3)$ equivariance.
+That is, under this perspective, there are no specific requirements or special treatments needed for the network design, as the $O(3)$ transformations applied to the crystal can be externally determined and simultaneously applied directly to the output.
+Building upon Proposition 3.4, the key challenge in achieving $O(3)$-equivariant tensor prediction lies in identifying the appropriate orbit canonicalization and rigid registration functions.
 3.3 Our Proposed Framework: GoeCTP
-In what follows, we will first introduce the core rotation and reflection (R&R) module of the proposed GoeCTP, focusing on how to obtain a standardized crystal position for a crystal with arbitrary spatial direction using polar decomposition.
+In what follows, we demonstrate that polar decomposition can effectively serve as both orbit canonicalization and rigid registration functions for crystal tensor prediction.
+Building on this foundation, we will first introduce the core rotation and reflection (R&R) module of the proposed GoeCTP, which is to obtain the canonical form for a crystal with arbitrary spatial orientations using polar decomposition.
 Then, we will describe how the input crystal data is processed and introduce the property prediction network of GoeCTP.
-Finally, we will explain how proposed GoeCTP achieves equivariant property predictions.
+Finally, we will explain how proposed GoeCTP achieves equivariant tensor predictions.
 An overview of the GoeCTP framework is illustrated in Fig.
 2.
+Theorem 3.5.
+(Polar Decomposition [9, 12, 16].) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\mathbf{H}$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{H}\in\mathbb{R}^{3\times 3}$ is a Hermitian positive semi-definite matrix.
+In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for rotating the crystal configuration to the canonical form.
+Therefore, throughout this work, we adopts the fractional coordinate system to represent the crystal represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$.
+Then building on Theorem 3.5 , we can define both orbit canonicalization and rigid registration functions for crystal tensor prediction as follows.
+Proposition 3.6.
+(Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix.
+In this perspective, $f_{p1}(\cdot)$ serves as the orbit canonicalization function, while $f_{p2}(\cdot)$ serves as the rigid registration function.
+The crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ is thus identified as the canonical form.
+The detailed proof of Proposition 3.6 is provided in Appendix A.7.
+Specifically, since the polar decomposition of a lattice matrix always exists and it is unique, any orthogonal matrix acting on the lattice matrix can be separated through polar decomposition, yielding a unique $\mathbf{H}$.
+Thus, $\mathbf{H}$ can serve as the canonical form.
+Moreover, the rationale of using polar decomposition is that it is a continuous canonicalization, which can provide improved robustness [6].
+For clarity, we provide details on this continuity in Appendix A.14, and further discussions on constructing alternative canonical forms based on polar decomposition in Appendix A.15, respectively.
+Building upon Proposition 3.6, we utilize polar decomposition to construct the proposed GoeCTP model.
 R&R Module.
-The primary function of the R&R Module is to rotate and reflect the crystal to a standardized crystal position in space.
-In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for the R&R module to perform rotation and reflection operations on the crystal data.
-Therefore, R&R module adopts the fractional coordinate system to represent the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$.
-This choice simplifies the implementation of the R&R module.
-Specifically, the application of $O(3)$-group transformations is achieved by applying polar decomposition as follows.
-Proposition 3.1
-(Polar Decomposition (Hall & Hall, 2013; Higham, 1986; Jiao et al., 2024).) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\exp(\mathbf{S})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{S}\in\mathbb{R}^{3\times 3}$ is a symmetric matrix and $\exp(\mathbf{S})=\sum_{n=0}^{\infty}\frac{\mathbf{S}^{n}}{n!}$ defines the exponential mapping of $\mathbf{S}$.
-Building on the above proposition, it is evident that for a crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the invertible lattice matrix $\mathbf{L}$ can be uniquely represented by a symmetric matrix $\mathbf{S}$.
-By defining $\mathbf{H}=\exp(\mathbf{S})$, the lattice matrix $\mathbf{L}$ can also be uniquely represented by a symmetric matrix denoted as $\mathbf{H}$, i.e.
-$\mathbf{L}=\mathbf{Q}\mathbf{H}$.
-Furthermore, $\mathbf{H}$ remains invariant with the $O(3)$ group transformation.
-Proof 3.1.1
-Given the lattice matrices $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$, where $\mathbf{Q^{\prime}}$ is an arbitrary $O(3)$ group transformation, applying polar decomposition on both $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$ yields the following expressions:
-$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (4)
-$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (5)
-By combining Eq. 4 and 5, we can have the following equation:
-$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (6)
-Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{T}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{T}\mathbf{Q^{\prime}}^{T}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix.
-According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$.
-If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition.
-Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation.
-In essence, regardless of the $O(3)$ group transformation applied to the crystal in space, the corresponding $\mathbf{H}$ associated with $\mathbf{L}$ remains invariant.
-Consequently, $\mathbf{H}$ serves as a fixed standardized crystal position for the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, allowing the crystal with different position to be consistently rotated to a same position through polar decomposition.
-The concept of the standardized crystal position can be formally stated as the following proposition :
-Proposition 3.2
-(Standardized Crystal Position.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$, the standardized crystal position $\mathbf{H}$ is obtained.
-When an arbitrary $O(3)$ group transformation $\mathbf{Q^{\prime}}$ is applied to a crystal structure, $\mathbf{H}$ remains unchanged.
-Therefore, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ formulated as:
-$f_{p}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H}),$ (7)
-where $f_{p}(\mathbf{M})$ represents the application of polar decomposition to the lattice matrix $\mathbf{L}$ of $\mathbf{M}$.
+The primary function of the R&R Module is to rotate the crystal configuration to its canonical form and obtain the corresponding transformation.
+Therefore, based on Proposition 3.6, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ to construct the rigid registration and the orbit canonicalization.
+Formally, the R&R Module can be expressed as:
+$f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H}),\text{ }f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q},$ (4)
+where $f_{p1}(\cdot)$ is the orbit canonicalization function, while $f_{p2}(\cdot)$ is the rigid registration function.
 As shown in Fig.
-2, the crystal $(\mathbf{A},\mathbf{F},\mathbf{H})$ is passed to the subsequent crystal graph construction module for further processing.
+2, the crystal configuration $(\mathbf{A},\mathbf{F},\mathbf{H})$ derived from orbit canonicalization is passed to the subsequent crystal graph construction module for further processing.
 The orthogonal matrix $\mathbf{Q}$ obtained during this decomposition is passed to the reverse R&R module, ensuring the equivariant transformation of the output tensor properties.
-Our proposed R&R module based on polar decomposition allows the input crystal data to be transformed into a standardized spatial position that is invariant under $O(3)$ space group transformations.
-With this particular module, the equivariance can be captured, meaning that the subsequent components of GeoCTP are no longer required to account for equivariance.
+Our proposed R&R module based on polar decomposition allows the input crystal data to be transformed into the canonical form that is invariant under $O(3)$ space group transformations.
+With this particular module, the equivariance can be captured, meaning that the subsequent components of GoeCTP are no longer required to account for equivariance.
 Crystal Graph Construction.
@@
 Here, we use the crystal graph construction from Yan et al.
-(2024a; b) to describe the structure and relationships within crystals.
-Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{P},\mathbf{H})$ as introduced in Sec.
-2.
-We assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$.
-$\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$.
-When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e.
-$d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$.
-Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$.
-Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al.
-(2024a); Yan et al.
-(2024a).
+[50, 51] to describe the structure and relationships within crystals.
+A detailed explanation of the crystal graph construction can be found in Appendix A.2.
+Upon construction, the edge features are denoted as $e_{ij}$, while the node features, representing atomic properties, are denoted as $\bm{a}_{i}$.
 Node and Edge Feature Embedding.
-Building on previous work (Xie & Grossman, 2018; Yan et al., 2024b; a), node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$.
+Building on previous work [47, 51, 50], node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$.
 Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$.
-The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding (Lin et al., 2023).
+The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding [21].
 Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
@@
 Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al.
-(2024a); Taniai et al.
-(2024); Yan et al.
-(2022); Lee et al.
-(2024); Wang et al.
-(2024a), among others.
-For better performance, we select eComformer (Yan et al., 2024a), which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module.
+[50], Taniai et al.
+[42], Yan et al.
+[49], Lee et al.
+[20], Wang et al.
+[44], among others.
+For better performance, we select eComformer [50], which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module.
 A detailed explanation of eComFormer can be found in Appendix A.2.
-Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows:
-$\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}.$ (8)
+Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows: $\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}$.
 Reverse R&R Module.
@@
 First, the Reverse R&R module transforms $\bm{G}^{\mathrm{final}}$ into a tensor output, as follows:
-$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}}),$ (9)
+$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}})$ (5)
 $\bm{\varepsilon}^{\mathrm{final}}=f_{rp}(\bm{\varepsilon},\mathbf{Q}),$ (6)
@@
 This conversion $f_{rp}(\cdot)$ for predicting the dielectric tensor can be expressed by:
-$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{T}.$ (11)
-For predicting the higher-order elastic tensor, the conversion process becomes more complex. Specifically, the conversion is expressed as follows:
-$C_{ijkl}^{\mathrm{final}}=\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{p=1}^{3}\sum_{q=1}^{3}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq},$ (12)
-where the subscripts denote the specific positions of the elements within the tensor data.
-4 Related work
-GNN-Based Methods.
-CGCNN is a pioneering GNN model specifically designed for handling crystal structures (Xie & Grossman, 2018).
-This model proposed to represent crystal structures as multi-edge crystal graphs.
-It was applied to predict various single-value properties such as formation energy and band gap.
-Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Das et al., 2023; Lin et al., 2023).
-These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors.
-Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction.
-In contrast, recent studies attempted to ensure equivariance through specialized network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023).
-These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties.
-Within these network architectures, many operations are required, such as tensor products and combining irreducible representations.
-These processes significantly increase computational costs, especially when handling higher-order data.
-Transformer-Based Methods.
-Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties.
-Matformer (Yan et al., 2022), one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells.
-This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures.
-Subsequently, more advanced Transformer-based approaches were proposed.
-Some methods, such as ComFormer (Yan et al., 2024a), CrystalFormer (Wang et al., 2024a), and CrystalFormer (Wang et al., 2024a), typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties.
-Furthermore, DOSTransformer (Lee et al., 2024) is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance.
-Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
+$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}.$ (7)
+For predicting the higher-order piezoelectric and elastic tensor, the conversion process becomes more complex, see Appendix A.8 for more details.
 4 Experiments
@@
 Datasets.
-In this work, we evaluate the performance of GoeCTP on two key tensor property prediction tasks: the second-order dielectric tensor and the fourth-order elastic tensor, respectively.
-The dataset for dielectric tensor is derived from the data processed by Yan et al.
-(2024b), sourced from the JARVIS-DFT database.
-Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/.
-The statistical details of the datasets are presented in Table 1.
-Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix.
-Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.2 for details) and then reconstruct the entire $3\times 3$ symmetric matrix.
-In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor (Itin & Hehl, 2013; Wen et al., 2024).
-Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.2 for details).
-Dataset Sample size Fnorm Mean Fnorm STD Unit
-Dielectric 4713 14.7 18.2 Unitless
-Elastic 25110 306.4 238.4 GPa
-Table 1: Dataset statistics.
-Baseline Methods.
-we selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e.
-MEGNET (Chen et al., 2019; Morita et al., 2020), EGTNN (Zhong et al., 2023), and GMTNet (Yan et al., 2024b), as baseline methods.
+In this work, we evaluate the performance of GoeCTP on three key tensor property prediction tasks: the second-order dielectric tensor, the third-order piezoelectric tensor, and the fourth-order elastic tensor, respectively.
+Further details of the dataset and experimental setup can be found in Appendix A.10.
+Baseline Methods. We selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET [1, 31], EGTNN [53], and GMTNet [51], as baseline methods.
 […]
-Experimental Settings.
-The experiments were performed using an NVIDIA GeForce RTX 3090 GPU.
-For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al.
-(2024b).
-For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio.
-During model training, Huber loss, AdamW (Loshchilov & Hutter, 2018), a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
-Since we observed some randomness in the results during the training process of the repeated dielectric dataset experiments, we conducted 5 repeated experiments for each method and used the average of the 5 experiment metrics as the final experimental results.
-The results of the 5 repeated experiments can be found in the Appendix 11.
 4.2 Experimental Results
@@
 This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet.
-Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024).
-Detailed results of these combinations are provided in Appendix 11.
+Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer [50] and CrystalFormer [42].
+Detailed results of these combinations are provided in Appendix A.11.
+In addition to these experimental results, we further investigate the impact of different canonical forms on tensor property prediction in crystalline materials.
+Detailed analyses can be found in Appendix A.16.
 MEGNET ETGNN GMTNet GoeCTP (Ours)
@@
 Table 1: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the dielectric dataset.
+Predicting Piezoelectric Tensors.
+The experimental results for the elastic tensor dataset are shown in Table 2.
+Similar to the experimental results on the dielectric tensors dataset, it achieves optimal results on all evaluation metrics.
+Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
+MEGNET ETGNN GMTNet GoeCTP (Ours)
+Fnorm $\downarrow$ 0.504 0.515 0.450 0.431
+EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9%
+EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9%
+EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7%
+Table 2: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the piezoelectric dataset.
 Predicting Elastic Tensors.
@@
 Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science.
-Further details on the results of GoeCTP’s combination with other models can be found in Appendix 11.
+Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
 MEGNET ETGNN GMTNet GoeCTP (Ours)
@@
 Table 3: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the elastic dataset.
+eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
+Dielectric: Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
+Total Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
+Piezoelectric: Fnorm $\downarrow$ 0.431 0.527 0.431 0.431
+Total Time (s) $\downarrow$ 37.83 37.23 38.01 37.71
+Elastic: Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
+Total Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
+Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric, piezoelectric, and elastic dataset.
+After training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets
 Verifying the $O(3)$ Equivariance.
@@
 Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set).
-All crystals in the original test set were adjusted to the standardized crystal position, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set.
-The method for generating the corresponding orthogonal matrices is from Heiberger (1978).
+All crystals in the original test set were adjusted to the canonical form, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set.
+The method for generating the corresponding orthogonal matrices is from Heiberger [10].
 We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
-The results on the dielectric tensor dataset are shown in Table 4.
-GoeCTP performed equally well on the augmented test set as on the original test set, indicating that it maintains strong $O(3)$ equivariance for tensor properties.
-In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45%, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties.
-The results on the elastic tensor dataset are shown in Table 5.
-Similar to the dielectric tensor dataset, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm dropping by nearly 29%.
-In contrast, GoeCTP’s performance remained unchanged, further demonstrating the effectiveness of the GoeCTP method.
-Additionally, runtime comparisons between GoeCTP and eComFormer on the test set showed that GoeCTP introduced no significant increase in runtime.
-This indicates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality without compromising efficiency.
-eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
-Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
-EwT 25% $\uparrow$ 83.2% 69.7% 83.2% 83.2%
-EwT 10% $\uparrow$ 56.8% 42.7% 56.8% 56.8%
-EwT 5% $\uparrow$ 35.5% 22.5% 35.5% 35.5%
-Total Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
-Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric dataset.
-eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
-Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
-EwT 25% $\uparrow$ 42.5% 25.9% 42.5% 42.5%
-EwT 10% $\uparrow$ 15.3% 2.2% 15.3% 15.3%
-EwT 5% $\uparrow$ 7.2% 0.2% 7.2% 7.2%
-Total Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
-Table 5: Ablation study for verifying the $O(3)$ equivariance with elastic dataset.
-Efficiency.
-The results presented in Table 6 and Table 7 illustrate the running times for GoeCTP and baseline methods.
-On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet.
-On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process.
-To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply.
-In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a multilayer perceptron (MLP) at the network’s output, ensuring both efficiency and scalability across different tensor orders.
+The results on the three tensor datasets are shown in Table 4, GoeCTP performed equally well on the augmented test set as on the original test set in all tensor datasets, indicating that it maintains strong $O(3)$ equivariance for tensor properties.
+In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45% in dielectric tensor dataset, nearly 22% in piezoelectric tensor dataset, and nearly 29% in elastic tensor dataset, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties.
 MEGNET ETGNN GMTNet GoeCTP (Ours)
 Total Time (s) $\downarrow$ 663 1325 1611 616
-Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048
+Time/batch (s) $\downarrow$ 0.043 0.085 0.103 0.039
 Table 5: Efficiency comparison on the dielectric dataset.
+Additionally, runtime comparisons between GoeCTP and eComFormer on the test set revealed that GoeCTP did not result in a significant increase in runtime.
+This demonstrates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality while maintaining efficiency.
+MEGNET ETGNN GMTNet GoeCTP (Ours)
+Total Time (s) $\downarrow$ 943 1220 5771 938
+Time/batch (s) $\downarrow$ 0.060 0.078 0.370 0.060
+Table 6: Efficiency comparison on the piezoelectric dataset.
+Efficiency.
+The results presented in Table 5, Table 6, and Table 7 illustrate the running times for GoeCTP and baseline methods.
+On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet.
+On the piezoelectric dataset, GoeCTP completed the entire training process with only 16.3% of the time spent compared to GMTNet.
+On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process.
 MEGNET ETGNN GMTNet GoeCTP (Ours)
 Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422
-Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189
+Time/batch (s) $\downarrow$ 0.036 0.056 $>$ 0.452 0.030
 Table 7: Efficiency comparison on the elastic dataset.
-6 Conclusion, limitations, and future works
-In this work, we propose a novel $O(3)$-equivariant framework GoeCTP for fast crystal tensor prediction, which is a plug-and-play framework that can be integrated with any existing single-value property prediction network, enabling them to predict tensor properties with almost no additional computational cost.
-Based on predicting invariant tensor properties from standardized crystal positions and using an external module to ensure tensor equivariance, this approach has achieved state-of-the-art performance and the highest efficiency in different widely used crystal tensor property benchmark tests.
-The limitations of our current GoeCTP include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module and, in theory, does not introduce additional performance enhancements.
-(2) Currently, GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types.
-In future work, we intend to explore the following directions: for (1), we aim to improve performance by integrating prior knowledge related to the independent components of tensor properties across different crystal systems in Appendix A.2; for (2), we plan to extend the framework to other domains, such as 3D point cloud research, broadening its applicability.
+To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply.
+In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a MLP at the network’s output, ensuring both efficiency and scalability across different tensor orders.
+5 Concluding Remarks
+In this work, we revisit crystal tensor property prediction through the lens of canonicalization.
+Built upon the notion of canonicalization, we proposed a novel $O(3)$-equivariant framework GoeCTP for fast and accurate crystal tensor prediction.
+Benefiting from the canonicalization mechanism, GoeCTP serves as a plug-and-play module that can be integrated with any existing scalar property prediction network, enabling it to predict tensor properties with negligible additional computational cost.
+The proposed approach achieves state-of-the-art predictive performance and superior efficiency across a variety of widely used benchmark datasets for crystal tensor properties.
+Except for the excel performance, our current GoeCTP has some limitations remain, which include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module.
+(2) GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types (See Appendix A.12 for more limitations details).
+In future study, we intend to improve performance by integrating prior knowledge related to the independent components of tensor properties across various crystal systems (see Appendix A.9).
+Further discussion on the utilization of space group constraints on tensor properties can be found in Appendix A.13.
 Appendix A Appendix
-A.1 Details of eComFormer
-The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks (Yan et al., 2024a).
+A.1 Related work
+GNN-Based Methods.
+CGCNN is a pioneering GNN model specifically designed for handling crystal structures [47].
+This model proposed to represent crystal structures as multi-edge crystal graphs.
+It was applied to predict various single-value properties such as formation energy and band gap.
+Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge [1, 26, 2, 3, 21].
+These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors.
+Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction.
+In contrast, recent studies attempted to ensure equivariance through specialized network architectures [28, 25, 11, 46, 34, 51, 53].
+These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties.
+Within these network architectures, many operations are required, such as tensor products and combining irreducible representations.
+These processes significantly increase computational costs, especially when handling higher-order data.
+Transformer-Based Methods.
+Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties.
+Matformer [49], one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells.
+This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures.
+Subsequently, more advanced Transformer-based approaches were proposed.
+Some methods, such as ComFormer [50], CrystalFormer [44], and CrystalFormer [42], typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties.
+Furthermore, DOSTransformer [20] is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance.
+Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
+Equivariant Learning.
+Recently, two general-purpose equivariant learning techniques for the arbitrary group have gained widespread attention: frame averaging [22, 36, 5] and canonicalization [27, 17, 6].
+Frame averaging is developed based on group averaging and involves averaging over smaller frames instead of the entire group.
+Canonicalization, on the other hand, is inherently related to frame averaging and can be viewed as a more general form of it.
+[27].
+In particular, Lin et al.
+[22] introduced minimal frame averaging, a method designed to achieve both efficient and exact equivalence in equivariant learning.
+This approach is primarily applicable to scenarios where the input and output spaces are identical, such as the n-body problem.
+It does not extend to cases involving different spaces, such as crystal tensor prediction.
+If this method were extended to tensor property prediction, it would represent a special case of Proposition 3.4, where both orbit canonicalization and rigid registration are achieved using QR decomposition.
+Within Proposition 3.4, however, different functions can be employed for orbit canonicalization and rigid registration.
+For instance, the Kabsch algorithm [18, 19] can be employed for rigid registration, while polar decomposition can be utilized for canonicalization.
+To the best of our knowledge, no previous work has extended these equivariant learning techniques to the prediction of crystal tensor properties.
+Thus, our method represents the first extension of these concepts to the crystal tensor property prediction.
+A.2 Details of Framework Architecture
+Crystal Graph Construction.
+In GoeCTP, we use the crystal graph construction from Yan et al.
+[50, 51] to describe the structure and relationships within crystals.
+Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{X},\mathbf{H})$ as introduced in Section 2.
+Then, we assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$.
+$\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$.
+When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e.
+$d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$.
+Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$.
+Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al.
+[44], Yan et al.
+[50].
+eComFormer.
+The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks [50].
 It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph.
 The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
-Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. (2024a).
+Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. [50].
 […]
 $\bm{\alpha}_{ij}=\frac{\bm{q}_{ij}\circ\bm{\xi}_{K}(\bm{k}_{ij})}{\sqrt{d_{\bm{q}_{ij}}}},\bm{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\bm{\alpha}_{ij}))\circ\bm{\xi}_{V}(\bm{\upsilon}_{ij}),$ (8)
-where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively, and the operators $\circ$ denote the Hadamard product, $\bm{\xi}_{msg}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$.
+where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively.
+The operators $\circ$ denote the Hadamard product, $\operatorname{BN}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$.
 Then, node feature $\bm{f}_{i}$ is updated as follows,
@@
 where $\bm{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
-The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers (Geiger & Smidt, 2022).
+The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers [7].
 It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features.
@@
 3.
-A.2 Tensor properties symmetry
-Tensor properties, such as dielectric tensors and elastic tensors, describe the material’s response to external physical fields (such as electric fields or mechanical stress).
+A.3 Proofs
+A.4 Proofs of Lemma A.1
+Lemma A.1.
+Given a orbit canonicalization function $C_{\mathbf{M}}(\cdot)$, for any prediction function ${f_{\theta}}(\cdot)$, it holds that ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ is $O(3)$-invariant.
+Proof.
+According to Definition 3.2, for any $g\in O(3)$ and $\mathbf{M}\in\mathcal{V}$, we have $C_{\mathbf{M}}(\mathbf{M})=\mathbf{M_{0}}$ and $C_{\mathbf{M}}(g\cdot\mathbf{M})=\mathbf{M_{0}}$. Thus,
+${f_{\theta}}(C_{\mathbf{M}}(g\cdot\mathbf{M}))$ $={f_{\theta}}(\mathbf{M_{0}})$ (13)
+$={f_{\theta}}(C_{\mathbf{M}}(\mathbf{M})),$
+which implies ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ is $O(3)$-invariant. ∎
+A.5 Proofs of Lemma A.2
+Lemma A.2.
+For a rigid registration function $R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})=g$, it is $O(3)$-equivariant with respect to the input $\mathbf{M}_{1}$.
+Proof.
+According to Definition 3.3, $\mathbf{M}_{1}=R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$. Thus, for any $g\in O(3)$, we have
+$R_{\mathbf{M}}(g\cdot\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$ $=g\cdot\mathbf{M}_{1}$ (14)
+$=g\cdot(R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2})$
+$=(g\cdot R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2}))\cdot\mathbf{M}_{2}.$
+Given $\mathbf{M}_{2}=(\mathbf{A}_{2},\mathbf{F}_{2},\mathbf{L}_{2})$, and noting that $g\cdot\mathbf{M}_{2}=(\mathbf{A}_{2},\mathbf{F}_{2},\mathbf{Q}\mathbf{L}_{2})$ (where $\mathbf{L}_{2}$ has full rank), we can write Eq.
+14 as
+$(R_{\mathbf{M}}(g\cdot\mathbf{M}_{1},\mathbf{M}_{2}))\mathbf{L}_{2}=(g\cdot R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2}))\mathbf{L}_{2}.$ (15)
+Therefore, $R_{\mathbf{M}}(g\cdot\mathbf{M}_{1},\mathbf{M}_{2})=g\cdot R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})$, which implies that $R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})$ is $O(3)$-equivariant.
+∎
+A.6 Proof of Proposition 3.4
+Proposition 3.4.
+($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
+Proof.
+According to Lemma A.2 and Lemma A.1, for any $g\in O(3)$ we have
+$h(g\cdot\mathbf{M})$ $=R_{\mathbf{M}}(g\cdot\mathbf{M},C_{\mathbf{M}}(g\cdot\mathbf{M}))\cdot f(C_{\mathbf{M}}(g\cdot\mathbf{M}))$ (16)
+$=R_{\mathbf{M}}(g\cdot\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$
+$=g\cdot R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$
+$=g\cdot h(\mathbf{M})$
+Thus, $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
+A.7 Proof of Proposition 3.6
+Proposition 3.6.
+(Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix.
+In this perspective, $f_{p1}(\cdot)$ serve as the orbit canonicalization function, while $f_{p2}(\cdot)$ serve as the rigid registration function.
+The matrix $\mathbf{H}$ is thus identified as the canonical form.
+Proof.
+Given the lattice matrices $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$, where $\mathbf{Q^{\prime}}$ is an arbitrary $O(3)$ group transformation, applying polar decomposition on both $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$ yields the following expressions:
+$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (17)
+$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (18)
+By combining Eq. 17 and 18, we can have the following equation:
+$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (19)
+Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{\top}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{\top}\mathbf{Q^{\prime}}^{\top}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix.
+According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$.
+If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition.
+Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation.
+Consequently, the crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ can serve as the canonical form.
+Performing polar decomposition on the lattice matrix to obtain $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$, i.e.
+$f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$, can thus be regarded as the orbit canonicalization function.
+Furthermore, based on Eq.
+17, performing polar decomposition on the lattice matrix to obtain $\mathbf{Q}$, i.e.
+$f_{p2}(\mathbf{M},\mathbf{M_{0}})=\mathbf{Q}$, can be regarded as rigid registration function, as shown below:
+$\mathbf{L}=\mathbf{Q}\mathbf{H}\quad\to\quad\mathbf{M}=\mathbf{Q}\mathbf{M_{0}}\quad\to\quad\mathbf{M}=f_{p2}(\mathbf{M},\mathbf{M_{0}})\cdot\mathbf{M_{0}}.$ (20)
+A.8 $O(3)$ equivariance for crystal tensor properties
+Due to the difference between the input space $\mathcal{V}$ and the output space $\mathcal{W}$, the $O(3)$ equivariance required for crystal tensor property prediction tasks differs from the $O(3)$ equivariance typically encountered in general molecular studies [13, 48, 40].
+In molecular studies, the input space $\mathcal{V}$ and the output space $\mathcal{W}$ are the same, meaning group representations in these spaces are identical, i.e.
+$\rho_{\mathcal{V}}(g)=\rho_{\mathcal{W}}(g)$.
+Specifically, for a function $f(\mathbf{X})$, if it is $O(3)$ equivariant, then it satisfies $f(\mathbf{Q}\mathbf{X})=\mathbf{Q}f(\mathbf{X})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an arbitrary orthogonal matrix, and $\mathbf{X}\in\mathbb{R}^{3\times N}$ represents the coordinate matrix of atoms in a molecule.
+In this context, the function $f$ can be seen as the model.
+However, in the prediction of crystal tensor properties, the requirements for $O(3)$ equivariance differ.
+In what follows, we will provide a detailed introduction to this difference.
+Dielectric tensor.
+For the dielectric tensor $\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$ and a function $f:(\mathbf{A},\mathbf{F},\mathbf{L})\to\bm{\varepsilon}$, if it is $O(3)$ equivariant, it must satisfy $f(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top}$.
+The reason for this difference lies in the physical nature of the dielectric tensor, which characterizes a material’s polarization response to an external electric field [51].
+Specifically, the dielectric tensor describes the relationship between the electric displacement $\mathbf{D}\in\mathbb{R}^{3}$ and the applied electric field $\mathbf{E}\in\mathbb{R}^{3}$ through the equation $\mathbf{D}=\bm{\varepsilon}\mathbf{E}$.
+When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal structure, the relationship $\mathbf{D^{\prime}}=\bm{\varepsilon}^{\prime}\mathbf{E^{\prime}}$ holds, where $\mathbf{D^{\prime}}=\mathbf{Q}\mathbf{D}$ and $\mathbf{E^{\prime}}=\mathbf{Q}\mathbf{E}$.
+Substituting these into the equation, we have:
+$\mathbf{Q}\mathbf{D}=\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E}\to\mathbf{D}=\mathbf{Q}^{\top}\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E},$ (21)
+which implies that the dielectric tensor transforms under the $O(3)$ group transformation as: $\bm{\varepsilon}^{\prime}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}$.
+This transformation principle extends similarly to other crystal tensors, as follows.
+Piezoelectric tensor.
+The piezoelectric tensor $\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ to the electric displacement field $\mathbf{D}\in\mathbb{R}^{3}$ within the material.
+Mathematically, this relationship is expressed as $\mathbf{D}_{i}=\sum_{jk}\mathbf{e}_{ijk}\bm{\epsilon}_{jk}$, with $i,j,k\in\{1,2,3\}$.
+When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and electric displacement field is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\mathbf{D}_{i}^{\prime}=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$.
+The relation is then reformulated as $\mathbf{D}_{i}^{\prime}=\sum_{jk}\mathbf{e}_{ijk}^{\prime}\bm{\epsilon}_{jk}^{\prime}$.
+Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$.
+Consequently, $\mathbf{D}_{i}^{\prime}$ can be represented as
+$\mathbf{D}_{i}^{\prime}$ $=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$ (22)
+$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}\bm{\epsilon}_{jk}$
+$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}(\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn})$
+$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{mn}\mathbf{e}_{\ell mn}(\sum_{jk}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}^{\prime}_{jk})\quad(exchange\,sign,m\leftrightarrow j,n\leftrightarrow k)$
+$=\sum_{jk}\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}\bm{\epsilon}_{jk}^{\prime}$
+Therefore, under the $O(3)$ group transformation $\mathbf{Q}$, the transformed piezoelectric tensor $\mathbf{e}_{ijk}^{\prime}$ is given by:
+$\mathbf{e}_{ijk}^{\prime}=\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}.$ (23)
+When employing GoeCTP to predict the piezoelectric tensor, Equation 7 becomes $\mathbf{e}_{ijk}^{\mathrm{final}}=\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}$.
+Elastic tensor.
+The elastic tensor $C\in\mathbb{R}^{3\times 3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ and the stress tensor $\sigma\in\mathbb{R}^{3\times 3}$ within the material.
+This relationship is expressed as $\bm{\sigma}_{ij}=\sum_{k\ell}C_{ijk\ell}\bm{\epsilon}_{k\ell}$, with $i,j,k,\ell\in\{1,2,3,4\}$.
+When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and stress tensor is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\bm{\sigma}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\sigma}_{mn}$.
+Under this transformation, the relationship becomes $\bm{\sigma}_{ij}^{\prime}=\sum_{k\ell}C_{ijk\ell}^{\prime}\bm{\epsilon}_{k\ell}^{\prime}$.
+Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$.
+Based on above equations, the transformed stress tensor $\bm{\sigma}_{ij}^{\prime}$ can be expressed as:
+$\bm{\sigma}_{ij}^{\prime}$ $=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\bm{\sigma}_{mn}$ (24)
+$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\bm{\epsilon}_{pq}$
+$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\sum_{k\ell}\mathbf{Q}_{kp}\mathbf{Q}_{\ell q}\bm{\epsilon}^{\prime}_{k\ell}$
+$=\sum_{k\ell}\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}\bm{\epsilon}^{\prime}_{k\ell}$
+Therefore, under the $O(3)$ group transformation $\mathbf{Q}$, $C_{ijkl}^{\prime}$ is represented as:
+$C_{ijkl}^{\prime}=\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}.$ (25)
+When GoeCTP is used to predict elastic tensor, Equation 7 becomes $C_{ijkl}^{\mathrm{final}}=\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}$.
+A.9 Tensor properties symmetry
+Tensor properties describe the material’s response to external physical fields (such as electric fields or mechanical stress).
 In materials with symmetry, this response must adhere to the symmetry requirements of the material.
 As demonstrated by Yan et al.
-(2024b), when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$,the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$.
-Therefore, the corresponding tensor properties also remain unchanged, i.e., $\bm{\varepsilon}=\mathbf{R}\bm{\varepsilon}\mathbf{R}^{T}$.
+[51], when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$.
+Therefore, the corresponding tensor properties also remain unchanged, i.e., $\bm{\varepsilon}=\mathbf{R}\bm{\varepsilon}\mathbf{R}^{\top}$.
 Thus, crystal symmetry imposes strict constraints on the components of the tensor, leading to the simplification or elimination of many components, reducing the number of independent components.
@@
 The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry.
-The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 8 (Mao et al., 2024).
+The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 8 [28].
 Voigt notation for elastic tensor.
@@
 The elastic tensor has a minimum of 3 and a maximum of 21 independent elements for various types of systems.
-Voigt notation is a compact way to represent these independent components of tensor properties (Itin & Hehl, 2013).
-According to the rules $11\to 1\mathrm{~{};~{}}22\to 2\mathrm{~{};~{}}33\to 3\mathrm{~{};~{}}23,~{}32\to 4\mathrm{~{};~{}}31,~{}13\to 5\mathrm{~{};~{}}12,~{}21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix (Wen et al., 2024; Ran et al., 2023):
+Voigt notation is a compact way to represent these independent components of tensor properties [14].
+According to the rules $11\to 1\mathrm{~{};~{}}22\to 2\mathrm{~{};~{}}33\to 3\mathrm{~{};~{}}23,~{}32\to 4\mathrm{~{};~{}}31,~{}13\to 5\mathrm{~{};~{}}12,~{}21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix [46, 37]:
 $C=\begin{pmatrix}C_{1111}&C_{1122}&C_{1133}&C_{1123}&C_{1131}&C_{1112}\\ C_{1122}&C_{2222}&C_{2233}&C_{2223}&C_{2231}&C_{2212}\\ C_{1133}&C_{2233}&C_{3333}&C_{3323}&C_{3331}&C_{3312}\\ C_{1123}&C_{2223}&C_{3323}&C_{2323}&C_{2331}&C_{2312}\\ C_{1131}&C_{2231}&C_{3331}&C_{2331}&C_{3131}&C_{3112}\\ C_{1112}&C_{2212}&C_{3312}&C_{2312}&C_{3112}&C_{1212}\end{pmatrix}\to\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}$
-The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to Wen et al. (2024); Ran et al. (2023)).
+The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to Wen et al. [46], Ran et al. [37]).
 Crystal system Number of independent elements Elastic tensor
@@
 Tetragonal 6 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&0&0&0\\ C_{12}&C_{11}&C_{13}&0&0&0\\ C_{13}&C_{13}&C_{33}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{44}&0\\ 0&0&0&0&0&C_{66}\end{pmatrix}\end{gathered}$
-Orthorhombic 9 $C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&0&0&0\\ C_{12}&C_{22}&C_{23}&0&0&0\\ C_{13}&C_{23}&C_{33}&0&0&0\\ 0&0&0&C_{44}&0&0\\ 0&0&0&0&C_{55}&0\\ 0&0&0&0&0&C_{66}\end{pmatrix}$
 Triclinic 21 $\begin{gathered}C=\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}\end{gathered}$
 Table 9: Number of independent components in the elastic tensor for different crystal systems.
-A.3 Experimental details and additional results
+Voigt notation for piezoelectric tensor.
+The number of independent components in the piezoelectric tensor for different crystal systems is shown in Tabel 10 (partial data shown; for more details, refer to De Jong et al.
+[4], Gorfman and Zhang [8]).
+Crystal system point groups Number of independent elements Piezoelectric tensor
+Trigonal 32 2 $\mathbf{e}=\begin{pmatrix}\mathbf{e}_{11}&-\mathbf{e}_{11}&0&\mathbf{e}_{14}&0&0\\ 0&0&0&0&-\mathbf{e}_{14}&-\mathbf{e}_{11}\\ 0&0&0&0&0&0\end{pmatrix}$
+Monoclinic 2 8 $\mathbf{e}=\begin{pmatrix}0&0&0&\mathbf{e}_{14}&0&\mathbf{e}_{16}\\ \mathbf{e}_{21}&\mathbf{e}_{22}&\mathbf{e}_{23}&0&e_{25}&0\\ 0&0&0&\mathbf{e}_{34}&0&\mathbf{e}_{36}\end{pmatrix}$
+Triclinic 1 18 $\mathbf{e}=\begin{pmatrix}\mathbf{e}_{11}&\mathbf{e}_{12}&\mathbf{e}_{13}&\mathbf{e}_{14}&\mathbf{e}_{15}&\mathbf{e}_{16}\\ \mathbf{e}_{21}&\mathbf{e}_{22}&\mathbf{e}_{23}&\mathbf{e}_{24}&\mathbf{e}_{25}&\mathbf{e}_{26}\\ \mathbf{e}_{31}&\mathbf{e}_{32}&\mathbf{e}_{33}&\mathbf{e}_{34}&\mathbf{e}_{35}&\mathbf{e}_{36}\end{pmatrix}$
+Table 10: Number of independent components in the piezoelectric tensor for different crystal systems.
+A.10 Experimental details
+Datasets details.
+The dataset for dielectric tensor and piezoelectric tensor is derived from the data processed by Yan et al.
+[51], sourced from the JARVIS-DFT database.
+Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/.
+The statistical details of the datasets are presented in Table 11.
+Additionally, in the dielectric tensor dataset, the dielectric tensor is a $3\times 3$ symmetric matrix.
+Therefore, during prediction, we predict 6 elements of the matrix (see Appendix A.9 for details) and then reconstruct the entire $3\times 3$ symmetric matrix.
+In the piezoelectric tensor dataset, the piezoelectric tensor is represented using Voigt notation as a $3\times 6$ matrix, rather than a $3\times 3\times 3$ third-order tensor (see Appendix A.9 for details).
+we only predict the $3\times 6$ matrix.
+In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor [14, 46].
+Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.9 for details).
+Dataset Sample size Fnorm Mean Fnorm STD Unit
+Dielectric 4713 14.7 18.2 Unitless
+Piezoelectric 4998 0.43 3.09 $\text{C}/\text{m}^{2}$
+Elastic 25110 306.4 238.4 GPa
+Table 11: Dataset statistics.
+Experimental Settings.
+The experiments were performed using a single NVIDIA GeForce RTX 3090 GPU.
+For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al.
+[51].
+For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio.
+During model training, Huber loss, AdamW [24], a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
 Hyperparameter settings of GoeCTP.
@@
 For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector.
-In the dielectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector.
-For the prediction module, eComFormer, in the dielectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer.
-For both dielectric and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64.
+In the dielectric and piezoelectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector.
+For the prediction module, eComFormer, in the dielectric and piezoelectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer.
+For both dielectric, piezoelectric, and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64.
 In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output.
+For the piezoelectric prediction task, the 18-dimensional vector features output by the prediction module are reconstructed into a $3\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
 For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
-1st 2nd 3rd 4th 5th mean
-MEGNET
-Fnorm $\downarrow$ 3.62 3.70 3.66 3.74 3.81 3.71
-EwT 25% $\uparrow$ 76.0% 77.1% 75.6% 73.7% 76.4% 75.8 %
-EwT 10% $\uparrow$ 41.2% 39.3% 38.2% 39.7% 36.3% 38.9%
-EwT 5% $\uparrow$ 17.6% 16.8% 18.5% 19.1% 17.8% 18.0%
-ETGNN
-Fnorm $\downarrow$ 3.42 3.48 3.33 3.35 3.41 3.40
-EwT 25% $\uparrow$ 82.1% 81.7% 83.4% 82.6% 83.2% 82.6 %
-EwT 10% $\uparrow$ 46.9% 47.8% 50.7% 49.4% 50.5% 49.1%
-EwT 5% $\uparrow$ 22.3% 26.5% 27.1% 25.9% 24.6% 25.3%
-GMTNet
-Fnorm $\downarrow$ 2.98 3.64 3.23 3.16 3.39 3.28
-EwT 25% $\uparrow$ 83.4% 81.3% 83.8% 84.5% 83.7% 83.3 %
-EwT 10% $\uparrow$ 54.9% 56.2% 55.8% 56.2% 57.1% 56.0%
-EwT 5% $\uparrow$ 31.2% 30.5% 28.8% 29.3% 32.7% 30.5%
-GoeCTP (Ours)
-Fnorm $\downarrow$ 3.33 3.29 3.36 3.25 2.92 3.23
-EwT 25% $\uparrow$ 82.5% 83.4% 82.1% 83.9% 84.2% 83.2 %
-EwT 10% $\uparrow$ 57.5% 53.9% 56.4% 56.7% 59.7% 56.8%
-EwT 5% $\uparrow$ 36.9% 31.6% 37.1% 34.2% 37.8% 35.5%
-Table 10: Comparison of 5 repeated experiments on dielectric dataset.
 Hyperparameter settings of GMTNet, ETGNN, and MEGNET.
 Following Yan et al.
-(2024b), we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks.
+[51], we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks.
 The same polynomial learning rate decay scheduler is used in all experiments.
+A.11 Additional results.
+We assessed the performance of GoeCTP when combined with iComFormer [50] and CrystalFormer [42] (denoted as GoeCTP (iCom.) for the combination) on both the dielectric, piezoelectric, and elastic tensor datasets.
+As presented in Table 12, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.).
+As shown in Table 13, similar to the dielectric tensor dataset, GoeCTP (eCom.) achieved the best prediction accuracy on the piezoelectric dataset, but GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.).
+In contrast, as shown in Table 14, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
 MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
@@
 Table 12: Additional comparison of performance metrics on dielectric dataset.
-Additional results.
-Given the relatively small size of the dielectric tensor dataset, along with other influencing factors, we observed some randomness in the results during repeated training runs, introducing noise into the outcome comparisons.
-To mitigate this effect and ensure the reliability of the results, we conducted five independent trials for each method, using the average of the performance metrics across these five trials as the final reported value.
-The detailed results from these repeated experiments are presented in Table 10.
-We assessed the performance of GoeCTP when combined with iComFormer (Yan et al., 2024a) and CrystalFormer (Taniai et al., 2024) (denoted as GoeCTP (iCom.) for the combination) on both the dielectric tensor and elastic tensor datasets.
-As presented in Table 11, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.).
-In contrast, as shown in Table 12, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
+MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
+Fnorm $\downarrow$ 0.504 0.515 0.450 0.431 0.448 0.457
+EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9% 45.5% 42.7%
+EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9% 44.9% 42.1%
+EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7% 38.3% 41.5%
+Total Time (s) $\downarrow$ 943 1220 5771 938 909 2787
+Time/batch (s) $\downarrow$ 0.074 0.095 0.45 0.073 0.071 0.218
+Table 13: Additional comparison of performance metrics on piezoelectric dataset.
 MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
@@
 Table 14: Additional comparison of performance metrics on elastic dataset.
+A.12 The limitations of our method and its broader impact
+Limitations.
+(1)As discussed in the main text, GoeCTP is a plug-and-play $O(3)$-equivariant framework designed to enhance the backbone network’s ability to achieve equivariant predictions.
+Consequently, the performance of our method is inherently dependent on the capabilities of the backbone network.
+If the backbone lacks sufficient predictive power, the combined framework may not fully surpass the state-of-the-art tensor prediction networks in all evaluation metrics.
+For instance, as shown in Table 12, when our framework is integrated with CrystalFormer [42], although the computational speed significantly exceeds that of GMTNet, the Fnorm, EwT 25%, and EwT 10% values are lower than those achieved by GMTNet.
+(2)As demonstrated in Higham [12] and Proposition 3.5, polar decomposition applied to a $3\times 3$ invertible matrix $\mathbf{L}$ produces a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$.
+For 3D crystal structures, the lattice matrix $\mathbf{L}$ is always full-rank (i.e.
+invertible), therefore, ensuring the applicability of our method to 3D crystal systems.
+However, for certain special cases, such as 2D crystals with single-layer structures [32, 39], the rank of the lattice matrix $\mathbf{L}$ may be less than 3.
+In these scenarios, directly applying polar decomposition may not yield a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$ and a unique canonical form $\mathbf{H}$.
+This limitation could cause our method to fail, as crystals with different spatial orientations cannot be consistently adjusted to a unique canonical form, thereby preventing the method from achieving the desired $O(3)$-equivariance.
+Broader impact.
+This work can be applied to screen existing crystal material databases for potential material candidates, facilitating the discovery of new materials with desirable tensor properties, such as high-dielectric materials.
+Therefore, our work has the potential to make a significant impact on materials science.
+A.13 How to further utilize the tensor properties symmetry
+In this section, we will use the dielectric tensor as an example to briefly discuss how to further utilize the symmetry of tensor properties.
+Utilizing symmetry constraints for zero elements.
+To demonstrate the role of symmetry constraints, we present an example of the GoeCTP prediction results in Table 15.
+Within a reasonable margin of error, the predictions align well with the expected constraints.
+Label Prediction Cubic dielectric tensor
+$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.252&0.016&0.008\\ 0.016&2.230&0.007\\ 0.008&0.007&2.262\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
+Table 15: An example of GoeCTP prediction results
+Crystal system Cubic Tetragonal Hexagonal-Trigonal Orthorhombic Monoclinic
+Success rate 88.3% 86.6% 84.5% 84.5% 75.7%
+Table 16: The GoeCTP results of predicting symmetry-constrained zero-valued dielectric tensor elements.
+For a dielectric tensor, we evaluate the prediction of zero elements by defining a threshold as 1% of the average value of non-zero elements in the labels.
+Predictions are deemed successful if they meet this threshold.
+The results are as shown in Table 16.
+It can be observed that our method predicts most zero elements accurately, though not flawlessly.
+Since our advantage lies in transferring equivariance through an external framework, there are no restrictions on the model itself.
+Therefore, to achieve a higher success rate for zero elements, we added a ReLU activation function to the output layer of the network to improve the success rate (this applies only to cases where tensor elements are greater than or equal to zero).
+The results after retraining GoeCTP are as shown in Table 17, Table 18, and Table 19.
+Our method predicts zero elements more accurately while maintaining overall performance.
+Label Prediction Cubic dielectric tensor
+$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.237&0.000&0.000\\ 0.000&2.283&0.000\\ 0.000&0.000&2.228\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
+Table 17: An example of GoeCTP (ReLU) prediction results
+Crystal system Cubic Tetragonal Hexagonal-Trigonal Orthorhombic Monoclinic
+Success rate 100% 100% 87.2% 100% 100%
+Table 18: The GoeCTP (ReLU) results of predicting symmetry-constrained zero-valued dielectric tensor elements.
+GoeCTP GoeCTP (ReLU)
+Fnorm $\downarrow$ 3.23 3.26
+EwT 25% $\uparrow$ 83.2% 82.6%
+EwT 10% $\uparrow$ 56.8% 58.4%
+EwT 5% $\uparrow$ 35.5% 36.3%
+Table 19: Performance comparisons between GoeCTP and GoeCTP (ReLU) on the dielectric dataset.
+This simple modification enables GoeCTP to more accurately predict the zero elements in dielectric tensors caused by the space group constraints.
+Utilizing symmetry constraints for non-zero elements.
+The symmetry constraints for zero elements serve as a foundational example.
+When the space group of the input crystal is known, prior knowledge (refer to Appendix A.9) can be employed to ensure that network outputs fully adhere to tensor property constraints.
+There are two specific methods to achieve this.
+In the first method, only the independent components of the tensor properties are predicted.
+For example, for a cubic crystal, we predict only one independent component and then use the mask from Table 8 to reconstruct it into a $3\times 3$ tensor).
+In Table 15, GoeCTP was modified to directly predict independent components.
+The new results are shown in Table 20, which fully satisfy the symmetry constraints.
+As shown in Table 21, the overall prediction performance indicates that directly predicting the independent components improves the model’s Fnorm performance on the dielectric dataset.
+Since this method directly predicts independent components, it is more suitable for cases like the dielectric tensor, where the independent components exhibit relatively simple structures.
+However, for more complex datasets, such as those involving piezoelectric and elastic tensors, the model’s performance using this method may degrade.
+In such scenarios, the second method provides a more viable alternative.
+Label Prediction Cubic dielectric tensor
+$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.357&0.000&0.000\\ 0.000&2.357&0.000\\ 0.000&0.000&2.357\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
+Table 20: An example of GoeCTP (predicting the independent components) prediction results
+GoeCTP GoeCTP (independent)
+Fnorm $\downarrow$ 3.23 3.11
+EwT 25% $\uparrow$ 83.2% 80.5%
+EwT 10% $\uparrow$ 56.8% 57.3%
+EwT 5% $\uparrow$ 35.5% 35.5%
+Table 21: Performance comparisons between GoeCTP and GoeCTP (predicting the independent components) on the dielectric dataset.
+In the second method, once the network training is completed, the network output is weighted using the standard tensor property format, referred to as a mask, as detailed in Table 8.
+The application of this method to the piezoelectric and elastic datasets is demonstrated in Tables 22 and 23.
+The results indicate that incorporating symmetry-based weighting effectively improves the performance of GoeCTP.
+GoeCTP GoeCTP (weighted)
+Fnorm $\downarrow$ 0.431 0.389
+EwT 25% $\uparrow$ 46.9% 48.5%
+EwT 10% $\uparrow$ 42.9% 46.7%
+EwT 5% $\uparrow$ 39.7% 46.1%
+Table 22: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the piezoelectric dataset.
+GoeCTP GoeCTP (weighted)
+Fnorm $\downarrow$ 107.11 106.24
+EwT 25% $\uparrow$ 42.5% 46.3%
+EwT 10% $\uparrow$ 15.3% 16.5%
+EwT 5% $\uparrow$ 7.2% 7.4%
+Table 23: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the elastic dataset.
+A.14 Analysis of discontinuity in canonicalization
+Previous studies [6] have shown that canonicalization can often introduce discontinuities into the model.
+To assess the impact of such discontinuities on our method, we analyze our proposed method from both theoretical and empirical perspectives.
+From theoretical perspective: The following proposition, presented in Appendix B of Dym et al. [6], characterizes the existence of a continuous canonicalization under orthogonal group actions:
+Proposition A.3.
+Consider the action of the orthogonal group $O(d)$ on $\mathbf{X}\in\mathbb{R}^{n\times d}$.
+There exists a continuous canonicalization for $O(d)$ when $n\leq d$.
+This continuous canonicalization can be defined as $y_{\mathbf{X}}=\left(\mathbf{X}^{\top}\mathbf{X}\right)^{1/2}$, where $\mathbf{X}^{\top}\mathbf{X}$ forms a positive semi-definite matrix and the square root is the standard square root of a positive semi-definite matrix, i.e.
+$\left(\mathbf{A}\right)^{1/2}$ is the unique positive semidefinite matrix $\mathbf{B}$ such that $\mathbf{B}^{2}=\mathbf{A}$.
+In fact, the polar decomposition adopted in this work can be regarded as a case of such continuous canonicalization.
+Proof.
+According to Proposition A.3, for a matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$, its continuous canonicalization norm is given by $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}$.
+In this work, the polar decomposition is applied to the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H},\mathbf{L}\in\mathbb{R}^{3\times 3}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{H}$ is the positive semi-definite Gram matrix (which serves as the canonicalization norm).
+Based on the properties of positive semi-definite matrices, our method is actually equivalent to the canonicalization approach described in Proposition A.3, i.e.
+$y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}=\left((\mathbf{Q}\mathbf{H})^{\top}\mathbf{Q}\mathbf{H}\right)^{1/2}=\left(\mathbf{H}^{\top}\mathbf{H}\right)^{1/2}=\mathbf{H}$.
+∎
+Therefore, our canonicalization via polar decomposition is consistent with the continuous canonicalization.
+From empirical perspective: To further demonstrate how this discontinuity affects the accuracy of tensor property predictions, we conducted a set of comparative experiments in which we replaced the polar decomposition in GoeCTP with QR decomposition [22] for evaluation.
+In this QR decomposition case, the lattice matrix is decomposed as $\mathbf{L}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{R}$ is an upper triangular matrix that serves as the canonicalization norm.
+However, $\mathbf{R}$ does not satisfy Proposition A.3, and therefore does not constitute a continuous canonicalization approach.
+To assess the sensitivity of each approach to small structural perturbations, we used the crystal $CdI_{2}$ (JARVIS_ID: JVASP-29631) as the model input, and then perturbed its lattice matrix (specifically, scaling down the first row of the lattice matrix) to observe the relative changes in the output (specifically the variations in the Fabenius norm of the output tensor).
+The detailed results are as follows:
+Perturbation Ratio 0% 5% 10% 15% 20% 25% 30% 35% 40%
+Output variation ratio (%): GoeCTP (QR) 0 3.38 6.63 11.29 21.83 39.37 67.81 124.79 191.28
+Output variation ratio (%): GoeCTP (polar) 0 2.53 1.23 2.03 4.84 12.72 24.88 36.34 45.67
+Table 24: Output variation ratio under different perturbation ratios on the dielectric dataset
+The results clearly show that the output of GoeCTP using polar decomposition exhibits significantly smaller relative changes compared to the QR-based variant, particularly under increasing levels of perturbation.
+This highlights the robustness of the continuous canonicalization afforded by polar decomposition, demonstrating its superiority in maintaining prediction stability and structural consistency under small input variations.
+A.15 How to design other canonical forms
+In fact, the canonical form obtained via polar decomposition provides a foundation from which infinitely many alternative canonical forms can be derived.
+Specifically, given a crystal with lattice matrix $\mathbf{L}$, polar decomposition yields $\mathbf{L}=\mathbf{Q}\mathbf{H}$.
+While $\mathbf{H}$ serves as one valid canonical representation, additional canonical forms can be constructed by applying orthogonal transformations to $\mathbf{H}$.
+For instance, given a fixed orthogonal matrix $\mathbf{Q}_{0}$ (which may be arbitrarily chosen) or a neural network $f_{Q}(\mathbf{H})=\mathbf{Q}_{\theta}$ that predicts an orthogonal matrix $\mathbf{Q}_{\theta}$, one can construct new canonical forms such as $\mathbf{Q}_{0}\mathbf{H}$ or $\mathbf{Q}_{\theta}\mathbf{H}$.
+This is because, for any given crystal with an arbitrary orientation, a fixed canonical form can be obtained according to a predefined computation rule, i.e.
+$\mathbf{L}=\mathbf{Q}\mathbf{H}\quad\to\quad\mathbf{L}=\mathbf{Q}\mathbf{Q}_{0}^{\top}\mathbf{Q}_{0}\mathbf{H}\quad\text{or}\quad\mathbf{L}=\mathbf{Q}\mathbf{Q}_{\theta}^{\top}\mathbf{Q}_{\theta}\mathbf{H}$.
+This formulation highlights the inherent flexibility in defining canonical forms.
+A.16 Impact of different canonical form on tensor property prediction
+As discussed in Appendix A.15, there exist infinitely many possible canonical forms.
+To investigate the impact of different canonical forms on tensor property prediction in crystalline materials, we first present the results comparing the performance of GoeCTP with QR decomposition [22] and GoeCTP without canonicalization.
+The corresponding results are summarized in the table below:
+GoeCTP (QR) GoeCTP (polar) GoeCTP (w/o QR or polar)
+Fnorm $\downarrow$ 3.20 3.23 3.60
+EwT 25% $\uparrow$ 83.5% 83.2% 80.6%
+EwT 10% $\uparrow$ 56.0% 56.8% 56.2%
+EwT 5% $\uparrow$ 32.4% 35.5% 32.6%
+Total Time (s) $\downarrow$ 639 616 613
+Table 25: Predictive performance comparisons between GoeCTP (QR), GoeCTP (polar), and GoeCTP (w/o QR or polar) on the dielectric dataset.
+According to the results, it is evident that incorporating a canonical form improves tensor prediction performance compared to omitting it.
+While different canonical forms lead to slight variations in performance, the overall impact remains relatively modest.
+One possible explanation is that canonical forms typically exhibit structured representations, such as the symmetric positive semi-definite matrix obtained via polar decomposition or the upper triangular matrix from QR decomposition, which may facilitate the model’s ability to capture relevant features.
+Furthermore, the use of a canonical form aligns all atoms within a crystal structure to a unified global coordinate system, thereby enabling equivariant tensor prediction.
+However, recent studies [43, 23, 15] have demonstrated that employing adaptive local frames, i.e.
+constructed based on individual atomic environments, can further enhance the accuracy of property predictions.
+Motivated by these insights, we extend the GoeCTP framework by incorporating local frames [43] (also referred to as local canonicalization).
+It is important to emphasize that GoeCTP can be seamlessly integrated with any prediction network, including those based on local frames, while preserving its equivariant prediction capability.
+The corresponding results are presented below.
+GoeCTP (local frame+polar) GoeCTP (polar)
+Fnorm $\downarrow$ 3.11 3.23
+EwT 25% $\uparrow$ 83.9% 83.2%
+EwT 10% $\uparrow$ 58.4% 56.8%
+EwT 5% $\uparrow$ 33.5% 35.5%
+Total Time (s) $\downarrow$ 1979 616
+Table 26: Predictive performance comparisons between GoeCTP (local frame+polar) and GoeCTP (polar) on the dielectric dataset.
+The results indicate that incorporating local frames can enhance predictive performance. However, this improvement comes at the expense of increased computational cost.

### v4

@@
 Abstract
-Predicting the tensor properties of crystalline materials is a fundamental task in materials science.
-Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance.
-Such equivariance constraint often requires specialized architecture designs to achieve effective predictions, inevitably introducing tremendous computational costs.
-Canonicalization, a classical technique for geometry, has recently been explored for efficient learning with symmetry.
-In this work, we revisit the problem of crystal tensor property prediction through the lens of canonicalization.
-Specifically, we demonstrate how polar decomposition, a simple yet efficient algebraic method, can serve as a form of canonicalization and be leveraged to ensure equivariant tensor property prediction.
-Building upon this insight, we propose a general $O(3)$-equivariant framework for fast and accurate crystal tensor property prediction, referred to as GoeCTP.
-By utilizing canonicalization, GoeCTP achieves high efficiency without requiring the explicit incorporation of equivariance constraints into the network architecture.
-Experimental results indicate that GoeCTP achieves the best prediction accuracy and runs up to 13$\times$ faster compared to existing state-of-the-art methods in benchmarking datasets, underscoring its effectiveness and efficiency.
-1 Introduction
-The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering [51].
+[…]
+Introduction
+The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering (yanspace).
 These tensor properties span various orders, such as dielectric tensor with two orders, piezoelectric tensor with three orders, and elastic tensor with four orders.
@@
 Thus far, several works have been dedicated to crystal tensor property prediction.
-One prominent category involves ab initio physical simulation techniques, such as density functional theory (DFT) [35].
+One prominent category involves ab initio physical simulation techniques, such as density functional theory (DFT) (petousis2016benchmarking).
 These classical simulation techniques can provide acceptable error margin for predicting various material properties.
-However, they necessitate extensive computational resources due to the complexity of handling large crystals with a vast number of atoms and electrons [51], hindering their applicability in practice.
+However, they necessitate extensive computational resources due to the complexity of handling large crystals with a vast number of atoms and electrons (yanspace), hindering their applicability in practice.
 Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction.
-These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) [1, 26, 2, 47] or transformers [50, 42, 49, 20, 44].
+These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) (chen2019graph; louis2020graph; choudhary2021atomistic; xie2018crystal) or transformers (yancomplete; taniaicrystalformer; yan2022periodic; lee2024density; wang2024conformal; ito2025rethinking).
 Most existing methods are designed for scalar property prediction, focusing on achieving $SO(3)$ invariance of the crystal structures.
 However, these scalar-property methods could not be adapted to predict tensor properties of crystalline materials, which exhibits significantly higher complexity.
-This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress [33, 38, 51]; their modeling necessitates preserving consistency with the crystal’s spatial orientation, exhibiting a unique tensor $O(3)$ equivariance [51].
-Therefore, a few recent studies attempt to ensure equivariance through specialized designs of network architectures [28, 25, 11, 46, 34, 51, 53].
-These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group.
-In this regard, numerous computationally intensive operations, such as tensor products and the merging of irreducible representations, are required.
-These operations incur large amounts of computation overhead, particularly when processing high-order data.
+This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress (nye1985physical; resta1994macroscopic; yanspace); their modeling necessitates preserving consistency with the crystal’s spatial orientation, exhibiting a unique tensor $O(3)$ equivariance (yanspace).
+Therefore, a few recent studies attempt to ensure equivariance through specialized designs of network architectures (mao2024dielectric; lou2024discovery; heilman2024equivariant; wen2024equivariant; pakornchote2023straintensornet; yanspace; zhong2023general).
+These methods generally integrate directional features and spherical harmonics to preserve equivariance.
+However, achieving this often requires computationally expensive operations such as tensor products, which introduce substantial overhead, especially when dealing with high-order tensor data.
 Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
-One approach that has show promising results in efficient learning with symmetry across various fields, including point clouds [22], n-body simulation [17], and antibodies generation [29], is the use of “canonicalization".
-Specifically, canonicalization maps a geometric data to an invariant representation [27, 6], referred to as the canonical form, and subsequently reconstructs an equivariant output from the canonical form without imposing any architectural constraints on the network.
+Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction.
+The visualization of the crystal structures on the left was generated using VESTA (momma2011vesta), while the visualization of the crystal tensor properties on the right follows the method of yanspace and VELAS (ran2023velas).
+One approach that has show promising results in efficient learning with symmetry across various fields, including point clouds (linequivariance), n-body simulation (kaba2023equivariance), and antibodies generation (martinkus2023abdiffuser), is the use of “canonicalization”.
+Specifically, canonicalization maps a geometric data to an invariant representation (ma2024canonicalization; dymequivariant), referred to as the canonical form, and subsequently reconstructs an equivariant output from the canonical form without imposing any architectural constraints on the backbone network.
 However, to date, it has not been explored for crystal tensor property prediction.
-In this work, we revisit the task of crystal tensor property prediction through the lens of canonicalization and introduce a novel canonicalization strategy tailored for this particular setting.
-Specifically, rather than embedding equivariance directly into model architecture, we propose a simple yet effective canonicalization module, termed R&R.
-In particular, our canonicalization is instantiated based on polar decomposition, a continuous mapping technique [6] that can provide enhanced robustness.
-During prediction, the R&R module applies rotations and reflections to transform the input crystal structure into its canonical form.
-The canonical form is then fed into a property prediction network to obtain the corresponding canonical tensor representation.
-Simultaneously, the orthogonal matrix derived from the R&R module is utilized to recover the equivariant output via the tensor transformation rule, enabling equivariant tensor property prediction without incurring additional computational overhead.
-We conducted experiments on dielectric, piezoelectric, and elastic tensor datasets, respectively, showcasing that the proposed method can achieve $O(3)$-equivariant tensor prediction while maintaining high efficiency.
-Compared to the previous state-of-the-art work Yan et al.
-[51], the GoeCTP method achieves higher prediction accuracy and runs by up to 13$\times$ faster in benchmarking datasets.
-2 Background
-2.1 Preliminaries
-The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell.
-An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters [49, 45, 16].
-General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
-Cartesian Coordinate System.
-A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{\top}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell.
-Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom.
-The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell.
-The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space.
-A complete crystal is therefore represented as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\bm{a}_{i}},\hat{\bm{x}}_{i})|\hat{\bm{x}}_{i}=\bm{x}_{i}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3},\hat{\bm{a}_{i}}=\bm{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$.
-In this representation, the integers $k_{i}$ and $l_{i}$ denote all possible atomic positions in the periodic lattice.
-Fractional Coordinate System.
-Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions.
-With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{\top}\in[0,1)^{3}$.
-The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$.
-Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
-Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA [30], Yan et al. [51], and VELAS [37].
-2.2 Problem statement
+[…]
+Figure 2: The Illustration of GoeCTP.
+To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary orientation, to the canonical form of this crystal.
+Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph.
+Subsequently, (4) the Prediction module leverages these embedded features to predict the canonical form of tensor properties corresponding to the canonical form of input crystal.
+Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the equivariance of the output tensor properties.
+Background
+Preliminaries
+Crystalline materials consist of a periodic arrangement of atoms in 3D space, characterized by a repeating unit known as the unit cell.
+A complete crystal can thus be described by the atomic types, atomic coordinates, and lattice parameters of a single unit cell (yan2022periodic; wang2024crystalline; jiaospace; huangcode; hua2025local).
+Mathematically, a crystal can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{n}]^{\top}\in\mathbb{R}^{n\times d_{a}}$ denotes the atomic features of $n$ atoms, $\mathbf{X}=[\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ contains their Cartesian coordinates, and $\mathbf{L}=[\boldsymbol{l}_{1},\boldsymbol{l}_{2},\boldsymbol{l}_{3}]\in\mathbb{R}^{3\times 3}$ is the lattice matrix composed of three basis vectors.
+The entire crystal can then be expressed as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\boldsymbol{a}_{i}},\hat{\boldsymbol{x}}_{i})|\hat{\boldsymbol{x}}_{i}=\boldsymbol{x}_{i}+k_{1}\boldsymbol{l}_{1}+k_{2}\boldsymbol{l}_{2}+k_{3}\boldsymbol{l}_{3},\hat{\boldsymbol{a}_{i}}=\boldsymbol{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$, which enumerates all atomic positions generated by lattice translations.
+Alternatively, using the lattice matrix $\mathbf{L}$ as the basis vectors leads to the fractional coordinate representation, where each atom is assigned a fractional coordinate $\boldsymbol{f_{i}}=[f_{i,1},f_{i,2},f_{i,3}]^{\top}\in[0,1)^{3}$ and its Cartesian coordinate is given by $\boldsymbol{x}_{i}=\sum_{j}f_{i,j}\boldsymbol{l}_{j}$.
+Accordingly, the crystal can also be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\boldsymbol{f}_{1},\cdots,\boldsymbol{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ denotes the fractional coordinates of all atoms in the unit cell.
+Problem statement
 Crystal Tensor Prediction.
@@
 Therefore, the crystal tensor property prediction task can be mathematically formulated as the following optimization problem:
-$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
+[…]
 where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset.
 In what follows, we will omit superscript $n$ for simplicity.
-As did in the literature [28, 46, 51], our objective is to estimate high-order tensor properties, including 2-order dielectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$), 3-order piezoelectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$), and 4-order elastic tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
+As previously proposed in the literature (mao2024dielectric; wen2024equivariant; yanspace), our objective is to estimate high-order tensor properties, including 2-order dielectric tensor (i.e., $\mathcal{Y}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\boldsymbol{\varepsilon}\in\mathbb{R}^{3\times 3}$), 3-order piezoelectric tensor (i.e., $\mathcal{Y}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$), and 4-order elastic tensor (i.e., $\mathcal{Y}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
 $\mathbf{O(3)}$ group.
@@
 $\mathbf{O(3)}$-Equivariance for Crystal Tensor Prediction.
-In the crystal tensor properties prediction task, due to the difference between $\rho_{\mathcal{V}}(g)$ and $\rho_{\mathcal{W}}(g)$, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies [13, 48, 52, 40].
-Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq.
-1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
+In the crystal tensor properties prediction task, due to the difference between $\rho_{\mathcal{V}}(g)$ and $\rho_{\mathcal{W}}(g)$, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies (hoogeboom2022equivariant; xuequivariant; zheng2024relaxing; song2024equivariant; aykent2025gotennet; cen2024high).
+Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\boldsymbol{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq.1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
 $f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top},$ (2)
-where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix [51].
-For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig.
-1 (for more equivariance details, see Appendix A.8).
-Figure 2: The Illustration of GoeCTP.
-To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary orientation, to the canonical form of this crystal.
-Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph.
-Subsequently, (4) the Prediction module leverages these embedded features to predict the canonical form of tensor properties corresponding to the canonical form of input crystal.
-Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the equivariance of the output tensor properties.
-3 Methodology
+where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an arbitrary orthogonal matrix (yanspace).
+For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig.1 (for more equivariance details, see Appendix A.4).
+Methodology
 In this section, we will first provide the motivation driving our new framework.
@@
 Building upon this foundational perspective, we demonstrate how polar decomposition can be employed as a canonicalization strategy and effectively applied to solve tensor prediction tasks with both effectiveness and efficiency.
-3.1 The Motivation of Our Framework
-To enforce $O(3)$ equivariance, as described in Eq.
-2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead.
-For example, existing work gather the rotational information from the neighboring nodes to the central node $i$ through the following computation formulated [51]:
-$\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),$ (3)
-where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ [7], and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics.
-To achieve $O(3)$ equivariance of the model and ensure high prediction accuracy for high-order tensor prediction tasks, such as elastic tensor prediction, the rotation order $\lambda$ is set to 3 [51], i.e., $\mathbf{Y}_{3}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{7}$.
-In contrast, for lower-order tensor prediction tasks, setting $\lambda$ to 1 is sufficient to achieve the model’s $O(3)$ equivariance, i.e., $\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{3}$.
-Compared to lower-order tensor prediction, the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ nearly doubles for high-order tensor predictions, leading to a significant increase in computational complexity.
-Consequently, this model architecture introduces tremendous computational costs for tensor property predictions.
-If we can achieve the model’s $O(3)$ equivariance and maintain the model performance for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, or even without relying on harmonic decomposition, computational costs could be drastically reduced.
-Inspired by the recent exploration of using canonicalization in efficient learning with symmetry, we aim to propose a principled approach based on canonicalization to enable $O(3)$-equivariant tensor predictions for crystal materials.
-3.2 $O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
-To mathematically describe $O(3)$-equivariant tensor prediction from the perspective of canonicalization, we first introduce and extend several fundamental concepts derived from canonicalization techniques [27, 17, 6].
-Definition 3.1.
-(Orbit.) The orbit of a crystal $\mathbf{M}$ is defined as $\text{Orbit}(\mathbf{M})=\{g\cdot\mathbf{M}\mid g\in O(3)\}$.
+The Motivation of Our Framework
+To enforce $O(3)$ equivariance as defined in Eq.
+2, existing crystal tensor prediction approaches typically rely on irreducible representations and Clebsch-Gordan (CG) tensor products.
+Although these formulations are theoretically rigorous and have shown strong empirical performance, they introduce substantial computational overhead due to the complexity of tensor operations.
+Recently, several studies have explored the scalarization trick to exploit high-degree steerable features, thereby eliminating the need for CG tensor products while maintaining both accuracy and efficiency (aykent2025gotennet; cen2024high; han2025survey).
+Similarly, if $O(3)$ tensor equivariance could be achieved for high-order tensor predictions without relying on tensor product operations, the computational cost could be significantly reduced.
+Motivated by recent progress in canonicalization-based symmetric learning, we propose a principled framework that leverages canonicalization to realize $O(3)$-equivariant tensor prediction for crystalline materials.
+$O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
+To mathematically describe $O(3)$-equivariant tensor prediction from the perspective of canonicalization, we first introduce and extend several fundamental concepts derived from canonicalization techniques (ma2024canonicalization; kaba2023equivariance; dymequivariant).
+Definition 1 (Orbit.).
+The orbit of a crystal $\mathbf{M}$ is defined as $\mathrm{Orbit}(\mathbf{M})=\{g\cdot\mathbf{M}\mid g\in O(3)\}$.
 In the context of crystal tensor prediction, all configurations of a crystal (i.e.
@@
 The formal process is outlined in the following definition.
-Definition 3.2.
-(Orbit Canonicalization.) A function denoted as $C_{\mathbf{M}}:\mathcal{V}\to\mathcal{V}$ is defined as an orbit canonicalization, if $\forall\mathbf{M}_{1}\in\text{Orbit}(\mathbf{M})$ and $\exists\mathbf{M}_{0}\in\text{Orbit}(\mathbf{M}),$ such that $C_{\mathbf{M}}(\mathbf{M}_{1})=\mathbf{M}_{0}$, where $\mathbf{M}_{0}$ is the canonical form of the orbit.
+Definition 2.
+(Orbit Canonicalization [^1] For simplicity, we refer to the procedure as canonicalization, though it may more precisely be regarded as a quasi-canonicalization that relaxes permutation invariance.
+) A function $C_{\mathbf{M}}:\mathcal{V}\to\mathcal{V}$ is defined as an orbit canonicalization if there exists a canonical form $\mathbf{M}_{0}\in\text{Orbit}(\mathbf{M})$ such that $C_{\mathbf{M}}(\mathbf{M}_{1})=\mathbf{M}_{0}$ for all $\mathbf{M}_{1}\in\text{Orbit}(\mathbf{M})$.
 Orbit canonicalization enables the transformation of a crystal with arbitrary orientations in the space into one with a specific, well-defined orientation.
@@
 To address this requirement, we introduce the concept of rigid registration.
-Rigid registration is a fundamental concept in fields such as computer vision and computational geometry [41].
+Rigid registration is a fundamental concept in fields such as computer vision and computational geometry (tam2012registration).
 It aims to find a transformation that matches the position and orientation of one object with the corresponding position and orientation of another object.
 This concept can be naturally adapted to the crystal, as follows.
-Definition 3.3.
+Definition 3.
 (Rigid Registration.) A function $R_{\mathbf{M}}:\mathcal{V}\times\mathcal{V}\to O(3)$ is a rigid registration if $\forall\mathbf{M}_{1},\mathbf{M}_{2}\in\text{Orbit}(\mathbf{M}),R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})=g$, such that $\mathbf{M}_{1}=g\cdot\mathbf{M}_{2}$.
-Using Definition 3.3, we can identify a transformation that matches two configurations of a crystal.
+Using Definition 3, we can identify a transformation that matches two configurations of a crystal.
 Specifically, this allows us to find specific group transformations that map all crystal configurations within the orbit to their canonical form.
-By leveraging Definition 3.2 and Definition 3.3 together, we derive a view for performing $O(3)$-equivariant tensor prediction through canonicalization, as described below.
-Proposition 3.4.
-($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function formulated as $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$-equivariant for tensor prediction.
-The proof of Proposition 3.4 is provided in Appendix A.6.
-Specifically, in this unified view of $O(3)$ equivariant tensor prediction, we leverage rigid registration, as defined in Definition 3.3, to find specific group transformations that map crystal configurations within the orbit to their canonical form.
-This canonical form is determined through orbit canonicalization, as described in Definition 3.2.
+By leveraging Definition 2 and Definition 3 together, we derive a view for performing $O(3)$-equivariant tensor prediction through canonicalization, as described below.
+Proposition 1.
+($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we define a new function $h(\mathbf{M})=R_{\mathbf{M}}\!\left(\mathbf{M},C_{\mathbf{M}}(\mathbf{M})\right)\!\cdot\!f\!\left(C_{\mathbf{M}}(\mathbf{M})\right)$, such that $h(\mathbf{M})$ is $O(3)$-equivariant for tensor prediction.
+The proof of Proposition 1 is provided in Appendix A.3.
+Specifically, in this unified view of $O(3)$ equivariant tensor prediction, we leverage rigid registration, as defined in Definition 3, to find specific group transformations that map crystal configurations within the orbit to their canonical form.
+This canonical form is determined through orbit canonicalization, as described in Definition 2.
 Subsequently, we can directly apply these transformations to the network’s output, thereby effectively achieving the $O(3)$ equivariance.
-That is, under this perspective, there are no specific requirements or special treatments needed for the network design, as the $O(3)$ transformations applied to the crystal can be externally determined and simultaneously applied directly to the output.
-Building upon Proposition 3.4, the key challenge in achieving $O(3)$-equivariant tensor prediction lies in identifying the appropriate orbit canonicalization and rigid registration functions.
-3.3 Our Proposed Framework: GoeCTP
+Under this perspective, our framework introduces no architectural constraints or special design requirements on the backbone network itself, since the $O(3)$ transformations of the crystal are externally determined through canonicalization and applied directly to the outputs.
+Building upon Proposition 1, the key challenge in achieving $O(3)$-equivariant tensor prediction lies in identifying the appropriate orbit canonicalization and rigid registration functions.
+Our Proposed Framework: GoeCTP
 In what follows, we demonstrate that polar decomposition can effectively serve as both orbit canonicalization and rigid registration functions for crystal tensor prediction.
@@
 Finally, we will explain how proposed GoeCTP achieves equivariant tensor predictions.
-An overview of the GoeCTP framework is illustrated in Fig.
-2.
-Theorem 3.5.
-(Polar Decomposition [9, 12, 16].) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\mathbf{H}$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{H}\in\mathbb{R}^{3\times 3}$ is a Hermitian positive semi-definite matrix.
+An overview of the GoeCTP framework is illustrated in Fig.2.
+Theorem 1.
+(Polar Decomposition (hall2013lie; higham1986computing; jiaospace).) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\mathbf{H}$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{H}\in\mathbb{R}^{3\times 3}$ is a symmetric positive semi-definite matrix.
 In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for rotating the crystal configuration to the canonical form.
-Therefore, throughout this work, we adopts the fractional coordinate system to represent the crystal represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$.
-Then building on Theorem 3.5 , we can define both orbit canonicalization and rigid registration functions for crystal tensor prediction as follows.
-Proposition 3.6.
+Therefore, throughout this work, we adopt the fractional coordinate system to represent the crystal represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$.
+Then building on Theorem 1, we can define both orbit canonicalization and rigid registration functions for crystal tensor prediction as follows.
+Proposition 2.
 (Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix.
@@
 The crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ is thus identified as the canonical form.
-The detailed proof of Proposition 3.6 is provided in Appendix A.7.
+The detailed proof of Proposition 2 is provided in Appendix A.3.
 Specifically, since the polar decomposition of a lattice matrix always exists and it is unique, any orthogonal matrix acting on the lattice matrix can be separated through polar decomposition, yielding a unique $\mathbf{H}$.
 Thus, $\mathbf{H}$ can serve as the canonical form.
-Moreover, the rationale of using polar decomposition is that it is a continuous canonicalization, which can provide improved robustness [6].
+Moreover, using polar decomposition is a continuous canonicalization, which can provide improved robustness (dymequivariant).
 For clarity, we provide details on this continuity in Appendix A.9, and further discussions on constructing alternative canonical forms based on polar decomposition in Appendix A.10, respectively.
-Building upon Proposition 3.6, we utilize polar decomposition to construct the proposed GoeCTP model.
+Building upon Proposition 2, we utilize polar decomposition to construct the proposed GoeCTP model.
 R&R Module.
 The primary function of the R&R Module is to rotate the crystal configuration to its canonical form and obtain the corresponding transformation.
-Therefore, based on Proposition 3.6, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ to construct the rigid registration and the orbit canonicalization.
+Therefore, based on Proposition 2, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ to construct the rigid registration and the orbit canonicalization.
 Formally, the R&R Module can be expressed as:
@@
 With this particular module, the equivariance can be captured, meaning that the subsequent components of GoeCTP are no longer required to account for equivariance.
-Crystal Graph Construction.
-To enable networks to handle such infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is typically necessary to employ graph construction methods that represent the interactions between infinite crystal structures and atoms using finite graph data.
-Here, we use the crystal graph construction from Yan et al.
-[50, 51] to describe the structure and relationships within crystals.
-A detailed explanation of the crystal graph construction can be found in Appendix A.2.
-Upon construction, the edge features are denoted as $e_{ij}$, while the node features, representing atomic properties, are denoted as $\bm{a}_{i}$.
-Node and Edge Feature Embedding.
-Building on previous work [47, 51, 50], node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$.
-Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$.
-The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding [21].
-Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
+Crystal Graph Preprocessing.
+To enable neural networks to effectively process infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is necessary to employ crystal graph construction and node and edge feature embedding techniques that represent atomic interactions within a finite graph representation.
+In this work, we adopt the crystal graph construction approach introduced by yancomplete; yanspace, where atoms are represented as nodes and interatomic interactions as edges.
+The node and edge features are subsequently embedded into vector representations $\boldsymbol{f}_{i}$ and $\boldsymbol{f}_{ij}^{e}$ to capture atomic and geometric information.
+A detailed description of the crystal graph construction procedure and the feature embedding scheme is provided in Appendix A.2.
 Prediction Module.
-Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al.
-[50], Taniai et al.
-[42], Yan et al.
-[49], Lee et al.
-[20], Wang et al.
-[44], among others.
-For better performance, we select eComformer [50], which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module.
+Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by yancomplete; taniaicrystalformer; yan2022periodic; lee2024density; wang2024conformal; ito2025rethinking, among others.
+In this work, we select eComFormer (yancomplete) as the Prediction module to illustrate the overall workflow of our proposed method.
 A detailed explanation of eComFormer can be found in Appendix A.2.
-Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows: $\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}$.
+Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows: $\boldsymbol{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\boldsymbol{f}_{i}^{\mathrm{final}}$.
 Reverse R&R Module.
 The primary function of the Reverse R&R module is to generate the equivariant tensor property predictions based on the crystal global features from the Prediction module.
-First, the Reverse R&R module transforms $\bm{G}^{\mathrm{final}}$ into a tensor output, as follows:
-$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}})$ (5)
-$\bm{\varepsilon}^{\mathrm{final}}=f_{rp}(\bm{\varepsilon},\mathbf{Q}),$ (6)
+First, the Reverse R&R module transforms $\boldsymbol{G}^{\mathrm{final}}$ into a tensor output, as follows:
+$\boldsymbol{\varepsilon}=f_{MLP}(\boldsymbol{G}^{\mathrm{final}})$ (4)
+$\boldsymbol{\varepsilon}^{\mathrm{final}}=f_{rp}(\boldsymbol{\varepsilon},\mathbf{Q}),$ (5)
 where $f_{MLP}(\cdot)$ represents a multilayer perceptron (MLP) and the operation reshaping dimension.
-Next, $f_{rp}(\cdot)$ utilizes the orthogonal matrix $\mathbf{Q}$ obtained from the R&R Module to convert the tensor output $\bm{\varepsilon}$ into its final equivariant form denoted as $\bm{\varepsilon}^{\mathrm{final}}$.
+Next, $f_{rp}(\cdot)$ utilizes the orthogonal matrix $\mathbf{Q}$ obtained from the R&R Module to convert the tensor output $\boldsymbol{\varepsilon}$ into its final equivariant form denoted as $\boldsymbol{\varepsilon}^{\mathrm{final}}$.
 This conversion $f_{rp}(\cdot)$ for predicting the dielectric tensor can be expressed by:
-$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}.$ (7)
+$\boldsymbol{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\boldsymbol{\varepsilon}\mathbf{Q}^{\top}.$ (6)
 For predicting the higher-order piezoelectric and elastic tensor, the conversion process becomes more complex, see Appendix A.4 for more details.
-4 Experiments
-4.1 Experimental setup
+Experiments
+Experimental setup
+Dataset Sample size Fnorm Mean Fnorm STD Unit
+Dielectric 4713 14.7 18.2 Unitless
+Piezoelectric 2701 0.79 4.03 $\text{C}/\text{m}^{2}$
+Elastic 25110 306.4 238.4 GPa
+Table 1: Dataset statistics.
 Datasets.
 In this work, we evaluate the performance of GoeCTP on three key tensor property prediction tasks: the second-order dielectric tensor, the third-order piezoelectric tensor, and the fourth-order elastic tensor, respectively.
+The datasets for dielectric tensor and piezoelectric tensor prediction are derived from the data curated by yanspace, which originates from the JARVIS-DFT database.
+Since nearly half of the crystal samples in the piezoelectric dataset have piezoelectric tensor labels that are entirely zero, training GoeCTP directly on this dataset can lead to severe overfitting.
+Therefore, we filtered out all crystal samples with zero-valued piezoelectric tensor labels from piezoelectric dataset.
+For the elastic tensor prediction task, the dataset is also obtained from the JARVIS-DFT database and is publicly accessible as dft_3d through the jarvis-tools package  [^2] https://pages.nist.gov/jarvis/databases/.
+The statistical details of the datasets are presented in Table 1.
 Further details of the dataset and experimental setup can be found in Appendix A.6.
-Baseline Methods. We selected several state-of-the-art methods in the field of crystal tensor property prediction, i.e. MEGNET [1, 31], EGTNN [53], and GMTNet [51], as baseline methods.
+Baseline Methods.
+We selected several state-of-the-art approaches in the field of crystal tensor property prediction, including EGTNN (zhong2023general) and GMTNet (yanspace), as baseline methods.
+Moreover, since GoeCTP is designed as a flexible framework, we further evaluated its performance when combined with different backbone models.
+In addition to the eComFormer (yancomplete) used in Figure 2, we also incorporated the recently proposed CrystalFramer (ito2025rethinking).
+Evaluation Metrics.
 […]
-4.2 Experimental Results
+Experimental Results
 Predicting Dielectric Tensors.
 The performance of various models in predicting the dielectric tensor is summarized in Table 2.
-While GoeCTP and GMTNet show close results in terms of Fnorm and EwT 25%, GoeCTP demonstrates higher values for EwT 10% and EwT 5%, with an improvement of 5% in EwT 5% compared to GMTNet.
-This indicates that GoeCTP delivers higher-quality predictions compared to GMTNet.
-Furthermore, as GoeCTP is designed to be a flexible framework, we also evaluated its combination with other models, such as iComFormer [50] and CrystalFormer [42].
-Detailed results of these combinations are provided in Appendix A.11.
+Since GoeCTP (C.
+Fra.) employs a more advanced backbone, its performance is notably superior to that of GoeCTP (C.
+eCom.).
+Both GoeCTP (eCom.) and GoeCTP (C.
+Fra.) outperform the baseline methods, demonstrating the effectiveness of the proposed GoeCTP framework.
+Comprehensive results combining GoeCTP with additional backbone architectures are provided in Appendix A.7.
 In addition to these experimental results, we further investigate the impact of different canonical forms on tensor property prediction in crystalline materials.
 Detailed analyses can be found in Appendix A.11.
-MEGNET ETGNN GMTNet GoeCTP (Ours)
-Fnorm $\downarrow$ 3.71 3.40 3.28 3.23
-EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2%
-EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8%
-EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5%
-Table 1: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the dielectric dataset.
+ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
+Fnorm $\downarrow$ 3.40 3.28 3.23 3.05
+EwT 25% $\uparrow$ 82.6% 83.3% 83.2% 86.4%
+EwT 10% $\uparrow$ 49.1% 56.0% 56.8% 62.6%
+EwT 5% $\uparrow$ 25.3% 30.5% 35.5% 43.5%
+Table 2: Comparison on the dielectric dataset.
+Lower Fnorm and higher EwT indicate better performance.
+GoeCTP (eCom.) denotes GoeCTP using eComFormer as the backbone, while GoeCTP (C.
+Fra.) refers to GoeCTP with CrystalFramer as the backbone.
 Predicting Piezoelectric Tensors.
-The experimental results for the elastic tensor dataset are shown in Table 2.
-Similar to the experimental results on the dielectric tensors dataset, it achieves optimal results on all evaluation metrics.
-Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
-MEGNET ETGNN GMTNet GoeCTP (Ours)
-Fnorm $\downarrow$ 0.504 0.515 0.450 0.431
-EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9%
-EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9%
-EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7%
-Table 2: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the piezoelectric dataset.
+The experimental results for the piezoelectric tensor dataset are shown in Table 3.
+Compared to the results on the dielectric dataset, predicting the piezoelectric tensor presents a greater challenge.
+As shown in Table 1, the overall mean Fnorm of the piezoelectric tensor is relatively small, which leads to lower overall EwT values.
+Consequently, achieving higher EwT requires more accurate predictions from the model.
+GoeCTP (C.
+Fra.) achieves relatively high EwT at both 5% and 10%, indicating that it attains high prediction accuracy for a subset of samples.
+Meanwhile, GMTNet and GoeCTP (eCom.) obtain better Fnorm results, suggesting that their overall prediction errors are comparatively smaller.
+ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
+Fnorm $\downarrow$ 0.873 0.752 0.778 0.827
+EwT 25% $\uparrow$ 0% 6.29% 2.59% 3.33%
+EwT 10% $\uparrow$ 0% 1.48% 1.14% 2.59%
+EwT 5% $\uparrow$ 0% 1.11% 0.04% 2.22%
+Table 3: Comparison on the piezoelectric dataset.
 Predicting Elastic Tensors.
-Experimental results on the elastic tensor dataset are presented in Table 3.
-GoeCTP demonstrated outstanding performance in predicting higher-order, complex tensors, outperforming all baseline methods across every evaluation metric.
-Notably, it achieved the lowest Fnorm of 107.11 and improved all EwT metrics by an average of 6% compared to GMTNet, emphasizing its general applicability in predicting diverse tensor properties in materials science.
-Further details on the results of GoeCTP’s combination with other models can be found in Appendix A.11.
-MEGNET ETGNN GMTNet GoeCTP (Ours)
-Fnorm $\downarrow$ 143.86 123.64 117.62 107.11
-EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5%
-EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3%
-EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2%
-Table 3: Comparison of performance metrics between MEGNET, ETGNN, GMTNet, and GoeCTP on the elastic dataset.
-eCom. (ori. data) eCom. (aug. data) GoeCTP (Ours) (ori. data) GoeCTP (Ours) (aug. data)
+Experimental results on the elastic tensor dataset are summarized in Table 4.
+Consistent with the observations on the dielectric dataset, both GoeCTP (eCom.) and GoeCTP (C.
+Fra.) achieved outstanding performance in predicting higher-order, complex tensors, surpassing all baseline methods across all evaluation metrics.
+Notably, GoeCTP (C.
+Fra.) delivered a further improvement over GoeCTP (eCom.).
+Additional results demonstrating GoeCTP combined with other backbone models are provided in Appendix A.7.
+ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
+Fnorm $\downarrow$ 123.64 117.62 107.11 95.98
+EwT 25% $\uparrow$ 32.0% 36.0% 42.5% 49.7%
+EwT 10% $\uparrow$ 3.8% 7.6% 15.3% 20.3%
+EwT 5% $\uparrow$ 0.5% 2.0% 7.2% 9.8%
+Table 4: Comparison on the elastic dataset.
+eCom. (ori.) eCom. (aug.) GoeCTP (ori.) GoeCTP (aug.)
 Dielectric: Fnorm $\downarrow$ 3.23 4.71 3.23 3.23
-Total Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
-Piezoelectric: Fnorm $\downarrow$ 0.431 0.527 0.431 0.431
-Total Time (s) $\downarrow$ 37.83 37.23 38.01 37.71
+Dielectric: Test Time (s) $\downarrow$ 26.03 26.01 26.23 26.18
+Piezoelectric: Fnorm $\downarrow$ 0.778 0.957 0.778 0.778
+Piezoelectric: Test Time (s) $\downarrow$ 5.21 5.40 5.53 6.05
 Elastic: Fnorm $\downarrow$ 107.11 138.45 107.11 107.11
-Total Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
-Table 4: Ablation study for verifying the $O(3)$ equivariance with dielectric, piezoelectric, and elastic dataset.
-After training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets
+Elastic: Test Time (s) $\downarrow$ 83.26 83.02 90.10 89.60
+Table 5: Ablation study for verifying the $O(3)$-equivariance using dielectric, piezoelectric, and elastic datasets.
+After training GoeCTP (eCom.), we extracted its prediction module (i.e., eComFormer) for comparative testing on two different test sets.
 Verifying the $O(3)$ Equivariance.
 To evaluate the effectiveness of GoeCTP, we conducted experiments to verify the $O(3)$ equivariance of tensor properties.
-Specifically, after training GoeCTP, we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set).
+Specifically, after training GoeCTP (eCom.), we extracted its Prediction module (i.e., eComFormer) for comparative testing on two different test sets (original test set and augmented test set).
 All crystals in the original test set were adjusted to the canonical form, while the augmented test set was generated by applying arbitrary $O(3)$ group transformations to all crystals in the original test set.
-The method for generating the corresponding orthogonal matrices is from Heiberger [10].
-We then evaluated both the Prediction module (eComFormer) and GoeCTP on these two datasets and compared the performance metrics.
+The method for generating the corresponding orthogonal matrices is from heiberger1978algorithm.
+We then evaluated both the Prediction module (eComFormer) and GoeCTP (eCom.) on these two datasets and compared the performance metrics.
 The results on the three tensor datasets are shown in Table 5, GoeCTP performed equally well on the augmented test set as on the original test set in all tensor datasets, indicating that it maintains strong $O(3)$ equivariance for tensor properties.
-In contrast, the performance of the eComFormer method significantly declined on the augmented test set, with Fnorm decreasing by nearly 45% in dielectric tensor dataset, nearly 22% in piezoelectric tensor dataset, and nearly 29% in elastic tensor dataset, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties.
-MEGNET ETGNN GMTNet GoeCTP (Ours)
-Total Time (s) $\downarrow$ 663 1325 1611 616
-Time/batch (s) $\downarrow$ 0.043 0.085 0.103 0.039
-Table 5: Efficiency comparison on the dielectric dataset.
+In contrast, the performance of the eComFormer method significantly declined on the augmented test set, demonstrating that it does not meet the $O(3)$ equivariance requirements for tensor properties.
 Additionally, runtime comparisons between GoeCTP and eComFormer on the test set revealed that GoeCTP did not result in a significant increase in runtime.
-This demonstrates that when integrated into single-value property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality while maintaining efficiency.
-MEGNET ETGNN GMTNet GoeCTP (Ours)
-Total Time (s) $\downarrow$ 943 1220 5771 938
-Time/batch (s) $\downarrow$ 0.060 0.078 0.370 0.060
-Table 6: Efficiency comparison on the piezoelectric dataset.
+This demonstrates that when integrated into scalar property prediction networks, GoeCTP incurs almost no additional computational cost, enhancing its practicality while maintaining efficiency.
+Dielectric: ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
+Total Time (s) $\downarrow$ 1325 1611 616 1976
+Time/epoch (s) $\downarrow$ 6.625 8.055 3.08 9.88
+Piezoelectric: ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
+Total Time (s) $\downarrow$ 736 7012 356 5768
+Time/epoch (s) $\downarrow$ 2.18 35.06 1.78 5.77
+Elastic: ETGNN GMTNet GoeCTP (eCom.) GoeCTP (C. Fra.)
+Total Time (s) $\downarrow$ 4448 >> 36000 2422 17197
+Time/epoch (s) $\downarrow$ 22.24 >> 180 12.11 34.394
+Table 6: Efficiency comparison.
 Efficiency.
-The results presented in Table 5, Table 6, and Table 7 illustrate the running times for GoeCTP and baseline methods.
-On the dielectric dataset, GoeCTP completed the entire training process with only 38.2% of the time spent compared to GMTNet.
-On the piezoelectric dataset, GoeCTP completed the entire training process with only 16.3% of the time spent compared to GMTNet.
-On the elastic dataset, GoeCTP required less than 7.0% of the time spent by GMTNet to complete the entire training process.
-MEGNET ETGNN GMTNet GoeCTP (Ours)
-Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422
-Time/batch (s) $\downarrow$ 0.036 0.056 $>$ 0.452 0.030
-Table 7: Efficiency comparison on the elastic dataset.
-To achieve $O(3)$ equivariance for tensor properties, GMTNet’s network architecture relies on irreducible representations and tensor operations, which considerably reduce computational efficiency, especially for higher-order tensor property (elastic tensor) prediction tasks where the time cost of GMTNet increases sharply.
-In contrast, GoeCTP requires no special architectural design for tensor property prediction tasks; it predicts tensor properties of varying orders using only a MLP at the network’s output, ensuring both efficiency and scalability across different tensor orders.
-5 Concluding Remarks
+Table 6 summarizes the training efficiency of GoeCTP and baseline methods.
+On the dielectric, piezoelectric, and elastic datasets, GoeCTP (eCom.) required only 38.2%, 5%, and 7% of GMTNet’s training time, respectively.
+Although GoeCTP (C.
+Fra.) achieved higher accuracy at the cost of longer training, it still maintained a efficiency advantage over GMTNet, particularly on high-order tensor datasets such as the elastic dataset.
+To achieve $O(3)$ equivariance for tensor properties, GMTNet employs a network architecture based on irreducible representations and tensor operations.
+Moreover, its final tensor output is obtained through gradient computation, which may significantly reduce computational efficiency.
+This inefficiency becomes particularly pronounced when computing gradients for high-order tensors such as elastic tensors.
+Beyond training costs, the gradient computation also may lead to slower inference.
+For instance, in the elastic tensor task, GMTNet requires more than 0.478 seconds per material, whereas GoeCTP (eCom.) requires only 0.036 seconds per material.
+In contrast, GoeCTP achieves equivariant tensor predictions without any specialized architectural constraints.
+It utilizes a simple multilayer perceptron (MLP) at the network output to predict tensor properties of various orders, ensuring high efficiency and scalability across different tensor orders.
+Conclusion
 In this work, we revisit crystal tensor property prediction through the lens of canonicalization.
@@
 Benefiting from the canonicalization mechanism, GoeCTP serves as a plug-and-play module that can be integrated with any existing scalar property prediction network, enabling it to predict tensor properties with negligible additional computational cost.
-The proposed approach achieves state-of-the-art predictive performance and superior efficiency across a variety of widely used benchmark datasets for crystal tensor properties.
-Except for the excel performance, our current GoeCTP has some limitations remain, which include (1) The performance of GoeCTP is inherently dependent on the performance of the Prediction module.
-(2) GoeCTP is specifically tailored for the prediction of tensor properties in crystalline materials and has not yet been adapted for other material types (See Appendix A.12 for more limitations details).
-In future study, we intend to improve performance by integrating prior knowledge related to the independent components of tensor properties across various crystal systems (see Appendix A.9).
-Further discussion on the utilization of space group constraints on tensor properties can be found in Appendix A.13.
+Except for the excel performance, our current GoeCTP has some limitations remain:
+  1. 1.
+    GoeCTP is a plug-and-play $O(3)$-equivariant framework designed to enhance the backbone network’s capability for equivariant prediction.
+Consequently, its overall performance inherently depends on the strength of the chosen backbone.
+For example, GoeCTP (eCom.) demonstrates higher efficiency, while GoeCTP (C.
+Fra.) achieves greater predictive accuracy.
+Striking an optimal balance between efficiency and accuracy remains a topic for further investigation.
+  2. 2.
+    As demonstrated in higham1986computing and Theorem 1, polar decomposition applied to an invertible matrix $\mathbf{L}$ produces a unique canonical form $\mathbf{H}$.
+For 3D crystal structures, the lattice matrix $\mathbf{L}$ is always full rank (i.e., invertible), ensuring the applicability of our approach to 3D crystal systems.
+However, in special cases such as 2D single-layer crystals (novoselov2005two; sherrell2022bright), the lattice matrix $\mathbf{L}$ may be rank-deficient.
+In these scenarios, polar decomposition may fail to yield a unique canonical form $\mathbf{H}$, making it impossible to consistently align crystals with different spatial orientations to a common canonical form, thereby hindering $O(3)$-equivariance.
+  3. 3.
+    The current framework underutilizes prior knowledge related to space group constraints and does not yet ensure that predicted tensor results strictly adhere to space group symmetry constraints.
+In future work, we plan to improve model performance by integrating priors concerning the independent tensor components across different crystal systems (see Appendix A.5).
+Further discussion on incorporating space group symmetry constraints into tensor property prediction is provided in Appendix A.8.
+Acknowledgments
+This work was partially supported by the Research Grants Counil (RGC) of the Hong Kong (HK) SAR (Grant No.
+15208725 and 15208222), the Young Scientists Fund of National Natural Science Foundation of China (NSFC) (Grant No.
+62206235), and the Hong Kong Polytechnic University (Grant No.
+A0046682 and P0057774).
 Appendix A Appendix
@@
 GNN-Based Methods.
-CGCNN is a pioneering GNN model specifically designed for handling crystal structures [47].
+CGCNN is a pioneering GNN model specifically designed for handling crystal structures (xie2018crystal).
 This model proposed to represent crystal structures as multi-edge crystal graphs.
-It was applied to predict various single-value properties such as formation energy and band gap.
-Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge [1, 26, 2, 3, 21].
-These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors.
+It was applied to predict various scalar properties such as formation energy and band gap.
+Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge (chen2019graph; louis2020graph; choudhary2021atomistic; das2023crysgnn; lin2023efficient).
+These GNN methods are primarily designed for scalar property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors.
 Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction.
-In contrast, recent studies attempted to ensure equivariance through specialized network architectures [28, 25, 11, 46, 34, 51, 53].
+In contrast, recent studies attempted to ensure equivariance through specialized network architectures (mao2024dielectric; lou2024discovery; heilman2024equivariant; wen2024equivariant; pakornchote2023straintensornet; yanspace; zhong2023general).
 These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties.
@@
 Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties.
-Matformer [49], one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells.
+Matformer (yan2022periodic), one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells.
 This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures.
 Subsequently, more advanced Transformer-based approaches were proposed.
-Some methods, such as ComFormer [50], CrystalFormer [44], and CrystalFormer [42], typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties.
-Furthermore, DOSTransformer [20] is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance.
+Some methods, such as ComFormer (yancomplete), CrystalFormer (wang2024conformal), CrystalFormer (taniaicrystalformer), and CrystalFramer (ito2025rethinking) typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of scalar properties.
+Furthermore, DOSTransformer (lee2024density) is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance.
 Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.
 Equivariant Learning.
-Recently, two general-purpose equivariant learning techniques for the arbitrary group have gained widespread attention: frame averaging [22, 36, 5] and canonicalization [27, 17, 6].
+Equivariance can be achieved through specialized architectural designs, such as incorporating high-degree steerable features (han2025survey).
+However, due to the structural constraints imposed by equivariant architectures, these models are often difficult to optimize (manolache2025learning; xie2025tale), and their complex designs may limit the universal approximation property (pacini2025universality; cenuniversally).
+Recently, two general-purpose techniques for achieving equivariant learning under arbitrary groups have gained widespread attention: frame averaging (linequivariance; punyframe; duval2023faenet) and canonicalization (ma2024canonicalization; kaba2023equivariance; dymequivariant).
+These methods represent promising alternatives to address the limitations of conventional equivariant models, as they impose no architectural constraints.
 Frame averaging is developed based on group averaging and involves averaging over smaller frames instead of the entire group.
 Canonicalization, on the other hand, is inherently related to frame averaging and can be viewed as a more general form of it.
-[27].
+(ma2024canonicalization).
 In particular, Lin et al.
-[22] introduced minimal frame averaging, a method designed to achieve both efficient and exact equivalence in equivariant learning.
+(linequivariance) introduced minimal frame averaging, a method designed to achieve both efficient and exact equivalence in equivariant learning.
 This approach is primarily applicable to scenarios where the input and output spaces are identical, such as the n-body problem.
 It does not extend to cases involving different spaces, such as crystal tensor prediction.
-If this method were extended to tensor property prediction, it would represent a special case of Proposition 3.4, where both orbit canonicalization and rigid registration are achieved using QR decomposition.
-Within Proposition 3.4, however, different functions can be employed for orbit canonicalization and rigid registration.
-For instance, the Kabsch algorithm [18, 19] can be employed for rigid registration, while polar decomposition can be utilized for canonicalization.
+If this method were extended to tensor property prediction, it would represent a special case of Proposition 1, where both orbit canonicalization and rigid registration are achieved using QR decomposition.
+Within Proposition 1.
+However, different functions can be employed for orbit canonicalization and rigid registration.
+For instance, the Kabsch algorithm (kabsch1976solution; lawrence2019purely) can be employed for rigid registration, while polar decomposition can be utilized for canonicalization.
 To the best of our knowledge, no previous work has extended these equivariant learning techniques to the prediction of crystal tensor properties.
@@
 A.2 Details of Framework Architecture
-Crystal Graph Construction.
-In GoeCTP, we use the crystal graph construction from Yan et al.
-[50, 51] to describe the structure and relationships within crystals.
-Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{X},\mathbf{H})$ as introduced in Section 2.
-Then, we assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$.
-$\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$.
-When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e.
-$d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$.
-Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$.
-Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al.
-[44], Yan et al.
-[50].
+[…]
+[…]
 eComFormer.
-The eComFormer has demonstrated strong performance across a range of single-value property prediction tasks [50].
+The eComFormer has demonstrated strong performance across a range of scalar property prediction tasks (yancomplete).
 It is built on an SO(3)-equivariant crystal graph representation, where the interatomic distance vectors are employed to represent the edge features of the graph.
 The model converts this crystal graph into embeddings and utilizes a transformer architecture, incorporating both a node-wise transformer layer and a node-wise equivariant updating layer, to extract rich geometric information during the message-passing process.
-Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from Yan et al. [50].
+Figure 3: The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer, adapted from yancomplete.
 […]
-$\bm{\alpha}_{ij}=\frac{\bm{q}_{ij}\circ\bm{\xi}_{K}(\bm{k}_{ij})}{\sqrt{d_{\bm{q}_{ij}}}},\bm{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\bm{\alpha}_{ij}))\circ\bm{\xi}_{V}(\bm{\upsilon}_{ij}),$ (8)
-where $\bm{\xi}_{K}$, $\bm{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively.
-The operators $\circ$ denote the Hadamard product, $\operatorname{BN}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\bm{q}_{ij}}}$ indicates the dimensionality of $\bm{q}_{ij}$.
-Then, node feature $\bm{f}_{i}$ is updated as follows,
-$\bm{m}s\bm{g}_{i}=\sum_{j\in\mathcal{N}_{i}}\bm{m}s\bm{g}_{ij},\bm{f}_{i}^{\mathrm{new}}=\bm{\xi}_{msg}(\bm{f}_{i}+\mathrm{BN}(\bm{m}s\bm{g}_{i})),$ (9)
-where $\bm{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
-The node-wise equivariant updating layer is designed to effectively capture geometric features by incorporating node feature $\mathbf{a}_{i}$ and edge feature $||\mathbf{e}_{ji}||_{2}$ as inputs and stacking two tensor product (TP) layers [7].
-It uses node feature $\bm{f}_{i}^{l}$ and equivalent vector feature $\mathbf{e}_{ji}$ embedded by corresponding spherical harmonics $\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})=c_{0}$,$\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})=c_{1}*\frac{\mathbf{e}_{ji}}{||\mathbf{e}_{ji}||_{2}}\in\mathbb{R}^{3}$ and $\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{5}$ to represent the input features.
-Gathering rotational information from neighboring nodes to the central node $i$, the first TP layer is shown as
-$\bm{f}_{i,0}^{l}=\bm{f}_{i}^{l^{\prime}}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})),\bm{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\bm{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),\lambda\in\{1,2\},$ (10)
-where $\bm{f}_{i}^{l^{\prime}}$ is linearly transformed from $\bm{f}_{i}^{l}$, $|\mathcal{N}_{i}|$ and $\mathbf{TP}_{\lambda}$ represent the number of neighbors of node $i$ and TP layers with rotation order $\lambda$ respectively.
+$\boldsymbol{\alpha}_{ij}=\frac{\boldsymbol{q}_{ij}\circ\boldsymbol{\xi}_{K}(\boldsymbol{k}_{ij})}{\sqrt{d_{\boldsymbol{q}_{ij}}}},\boldsymbol{msg}_{ij}=\operatorname{sigmoid}(\operatorname{BN}(\boldsymbol{\alpha}_{ij}))\circ\boldsymbol{\xi}_{V}(\boldsymbol{\upsilon}_{ij}),$ (7)
+where $\boldsymbol{\xi}_{K}$, $\boldsymbol{\xi}_{V}$ represent nonlinear transformations applied to key and value features, respectively.
+The operators $\circ$ denote the Hadamard product, $\operatorname{BN}(\cdot)$ refers to the batch normalization layer, and $\sqrt{d_{\boldsymbol{q}_{ij}}}$ indicates the dimensionality of $\boldsymbol{q}_{ij}$.
+Then, node feature $\boldsymbol{f}_{i}$ is updated as follows,
+$\boldsymbol{m}s\boldsymbol{g}_{i}=\sum_{j\in\mathcal{N}_{i}}\boldsymbol{m}s\boldsymbol{g}_{ij},\boldsymbol{f}_{i}^{\mathrm{new}}=\boldsymbol{\xi}_{msg}(\boldsymbol{f}_{i}+\mathrm{BN}(\boldsymbol{m}s\boldsymbol{g}_{i})),$ (8)
+where $\boldsymbol{\xi}_{msg}(\cdot)$ denoting the softplus activation function.
+[…]
+$\boldsymbol{f}_{i,0}^{l}=\boldsymbol{f}_{i}^{l^{\prime}}+\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j}^{l^{\prime}},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji})),\boldsymbol{f}_{i,\lambda}^{l}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{\lambda}(\boldsymbol{f}_{j}^{l^{\prime}},\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})),\lambda\in\{1,2\},$ (9)
+where $\boldsymbol{f}_{i}^{l^{\prime}}$ is linearly transformed from $\boldsymbol{f}_{i}^{l}$, $|\mathcal{N}_{i}|$ and $\mathbf{TP}_{\lambda}$ represent the number of neighbors of node $i$ and TP layers with rotation order $\lambda$ respectively.
 Then, to represent the invariant node features, the second TP layer is written as
-$\bm{f}_{i}^{l*}=\frac{1}{|\mathcal{N}_{i}|}(\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,0}^{l},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,1}^{l},\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\bm{f}_{j,2}^{l},\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji}))),$ (11)
-stacking the two tensor product layers together using both linear and nonlinear transformations, the output $\bm{f}_{i,updated}^{l}$ is combined as
-$\bm{f}_{i,updated}^{l}=\bm{\sigma}_{\mathrm{equi}}(\mathrm{BN}(\bm{f}_{i}^{l*}))+\mathrm{LN}_{\mathrm{equi}}(\bm{f}_{i}^{l}),$ (12)
-with $\bm{\sigma}_{\mathrm{equi}}$ denoting a nonlinear transformation made up of two softplus layers with a linear layer positioned between them.
+$\boldsymbol{f}_{i}^{l*}=\frac{1}{|\mathcal{N}_{i}|}(\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j,0}^{l},\mathbf{Y}_{0}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j,1}^{l},\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji}))+\sum_{j\in\mathcal{N}_{i}}\mathbf{TP}_{0}(\boldsymbol{f}_{j,2}^{l},\mathbf{Y}_{2}(\hat{\mathbf{e}}_{ji}))),$ (10)
+stacking the two tensor product layers together using both linear and nonlinear transformations, the output $\boldsymbol{f}_{i,updated}^{l}$ is combined as
+$\boldsymbol{f}_{i,updated}^{l}=\boldsymbol{\sigma}_{\mathrm{equi}}(\mathrm{BN}(\boldsymbol{f}_{i}^{l*}))+\mathrm{LN}_{\mathrm{equi}}(\boldsymbol{f}_{i}^{l}),$ (11)
+with $\boldsymbol{\sigma}_{\mathrm{equi}}$ denoting a nonlinear transformation made up of two softplus layers with a linear layer positioned between them.
 The detailed architectures of the node-wise transformer layer and node-wise equivariant updating layer are shown in Fig.
@@
 A.3 Proofs
-A.4 Proofs of Lemma A.1
-Lemma A.1.
+Proofs of Lemma 1
+Lemma 1.
 Given a orbit canonicalization function $C_{\mathbf{M}}(\cdot)$, for any prediction function ${f_{\theta}}(\cdot)$, it holds that ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ is $O(3)$-invariant.
 Proof.
-According to Definition 3.2, for any $g\in O(3)$ and $\mathbf{M}\in\mathcal{V}$, we have $C_{\mathbf{M}}(\mathbf{M})=\mathbf{M_{0}}$ and $C_{\mathbf{M}}(g\cdot\mathbf{M})=\mathbf{M_{0}}$. Thus,
+According to Definition 2, for any $g\in O(3)$ and $\mathbf{M}\in\mathcal{V}$, we have $C_{\mathbf{M}}(\mathbf{M})=\mathbf{M_{0}}$ and $C_{\mathbf{M}}(g\cdot\mathbf{M})=\mathbf{M_{0}}$. Thus,
 ${f_{\theta}}(C_{\mathbf{M}}(g\cdot\mathbf{M}))$ $={f_{\theta}}(\mathbf{M_{0}})$ (12)
@@
 which implies ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ is $O(3)$-invariant. ∎
-A.5 Proofs of Lemma A.2
-Lemma A.2.
+Proofs of Lemma 2
+Lemma 2.
 For a rigid registration function $R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})=g$, it is $O(3)$-equivariant with respect to the input $\mathbf{M}_{1}$.
 Proof.
-According to Definition 3.3, $\mathbf{M}_{1}=R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$. Thus, for any $g\in O(3)$, we have
+According to Definition 3, $\mathbf{M}_{1}=R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$. Thus, for any $g\in O(3)$, we have
 $R_{\mathbf{M}}(g\cdot\mathbf{M}_{1},\mathbf{M}_{2})\cdot\mathbf{M}_{2}$ $=g\cdot\mathbf{M}_{1}$ (13)
@@
 ∎
-A.6 Proof of Proposition 3.4
-Proposition 3.4.
+Proof of Proposition 1
+Proposition 1.
 ($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
 Proof.
-According to Lemma A.2 and Lemma A.1, for any $g\in O(3)$ we have
+According to Lemma 2 and Lemma 1, for any $g\in O(3)$ we have
 $h(g\cdot\mathbf{M})$ $=R_{\mathbf{M}}(g\cdot\mathbf{M},C_{\mathbf{M}}(g\cdot\mathbf{M}))\cdot f(C_{\mathbf{M}}(g\cdot\mathbf{M}))$ (15)
@@
 Thus, $h(\mathbf{M})$ is $O(3)$ equivariant for tensor prediction.
-A.7 Proof of Proposition 3.6
-Proposition 3.6.
+Proof of Proposition 2
+Proposition 2.
 (Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix.
@@
 Proof.
-Given the lattice matrices $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$, where $\mathbf{Q^{\prime}}$ is an arbitrary $O(3)$ group transformation, applying polar decomposition on both $\mathbf{L}$ and $\mathbf{Q^{\prime}}\mathbf{L}$ yields the following expressions:
-$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (17)
-$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (18)
-By combining Eq. 17 and 18, we can have the following equation:
-$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (19)
-Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{\top}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{\top}\mathbf{Q^{\prime}}^{\top}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix.
-According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$.
-If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition.
-Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation.
-Consequently, the crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ can serve as the canonical form.
-Performing polar decomposition on the lattice matrix to obtain $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$, i.e.
-$f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$, can thus be regarded as the orbit canonicalization function.
-Furthermore, based on Eq.
-17, performing polar decomposition on the lattice matrix to obtain $\mathbf{Q}$, i.e.
+Let $\mathbf{L}\in\mathbb{R}^{3\times 3}$ be an invertible lattice matrix.
+By the polar decomposition, there exist a unique orthogonal $\mathbf{Q}\in O(3)$ and a unique symmetric positive-definite $\mathbf{H}\in\mathbb{R}^{3\times 3}$ such that
+$\mathbf{L}=\mathbf{Q}\mathbf{H},\qquad\big(\mathbf{L}^{\top}\mathbf{L}\big)^{1/2}=\big(\mathbf{H}^{\top}\mathbf{H}\big)^{1/2}=\big(\mathbf{H}\mathbf{H}\big)^{1/2}=\mathbf{H}.$ (16)
+Consider an arbitrary $O(3)$ group transformation $\mathbf{Q^{\prime}}\in O(3)$ acting on the lattice, i.e., $\mathbf{L}\mapsto\mathbf{Q^{\prime}}\mathbf{L}$.
+Apply polar decomposition to $\mathbf{Q^{\prime}}\mathbf{L}$:
+$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q}_{1}\mathbf{H}_{1},\qquad\mathbf{H}_{1}=\big((\mathbf{Q^{\prime}}\mathbf{L})^{\top}(\mathbf{Q^{\prime}}\mathbf{L})\big)^{1/2}.$ (17)
+Since $(\mathbf{Q^{\prime}})^{\top}\mathbf{Q^{\prime}}=\mathbf{I}$, we have
+$\mathbf{H}_{1}=\big(\mathbf{L}^{\top}\mathbf{L}\big)^{1/2}=\mathbf{H}.$ (18)
+Therefore, $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation.
+Because $\mathbf{A}$ and $\mathbf{F}$ are unchanged by $O(3)$ group transformation and $\mathbf{H}$ is invariant by (18), the triplet
+$\mathbf{M}_{0}\coloneqq f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ (19)
+is identical for all representatives of the $O(3)$-orbit of $\mathbf{M}$. Therefore, $f_{p1}(\cdot)$ serves as an orbit canonicalization map and the matrix $\mathbf{H}$ is the canonical form.
+Furthermore, based on $\mathbf{L}=\mathbf{Q}\mathbf{H}$, performing polar decomposition on the lattice matrix to obtain $\mathbf{Q}$, i.e.
 $f_{p2}(\mathbf{M},\mathbf{M_{0}})=\mathbf{Q}$, can be regarded as rigid registration function, as shown below:
 $\mathbf{L}=\mathbf{Q}\mathbf{H}\quad\to\quad\mathbf{M}=\mathbf{Q}\mathbf{M_{0}}\quad\to\quad\mathbf{M}=f_{p2}(\mathbf{M},\mathbf{M_{0}})\cdot\mathbf{M_{0}}.$ (20)
-A.8 $O(3)$ equivariance for crystal tensor properties
-Due to the difference between the input space $\mathcal{V}$ and the output space $\mathcal{W}$, the $O(3)$ equivariance required for crystal tensor property prediction tasks differs from the $O(3)$ equivariance typically encountered in general molecular studies [13, 48, 40].
+A.4 $O(3)$ equivariance for crystal tensor properties
+Due to the difference between the input space $\mathcal{V}$ and the output space $\mathcal{W}$, the $O(3)$ equivariance required for crystal tensor property prediction tasks differs from the $O(3)$ equivariance typically encountered in general molecular studies (hoogeboom2022equivariant; xuequivariant; song2024equivariant).
 In molecular studies, the input space $\mathcal{V}$ and the output space $\mathcal{W}$ are the same, meaning group representations in these spaces are identical, i.e.
@@
 Dielectric tensor.
-For the dielectric tensor $\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$ and a function $f:(\mathbf{A},\mathbf{F},\mathbf{L})\to\bm{\varepsilon}$, if it is $O(3)$ equivariant, it must satisfy $f(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top}$.
-The reason for this difference lies in the physical nature of the dielectric tensor, which characterizes a material’s polarization response to an external electric field [51].
-Specifically, the dielectric tensor describes the relationship between the electric displacement $\mathbf{D}\in\mathbb{R}^{3}$ and the applied electric field $\mathbf{E}\in\mathbb{R}^{3}$ through the equation $\mathbf{D}=\bm{\varepsilon}\mathbf{E}$.
-When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal structure, the relationship $\mathbf{D^{\prime}}=\bm{\varepsilon}^{\prime}\mathbf{E^{\prime}}$ holds, where $\mathbf{D^{\prime}}=\mathbf{Q}\mathbf{D}$ and $\mathbf{E^{\prime}}=\mathbf{Q}\mathbf{E}$.
+For the dielectric tensor $\boldsymbol{\varepsilon}\in\mathbb{R}^{3\times 3}$ and a function $f:(\mathbf{A},\mathbf{F},\mathbf{L})\to\boldsymbol{\varepsilon}$, if it is $O(3)$ equivariant, it must satisfy $f(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top}$.
+The reason for this difference lies in the physical nature of the dielectric tensor, which characterizes a material’s polarization response to an external electric field (yanspace).
+Specifically, the dielectric tensor describes the relationship between the electric displacement $\mathbf{D}\in\mathbb{R}^{3}$ and the applied electric field $\mathbf{E}\in\mathbb{R}^{3}$ through the equation $\mathbf{D}=\boldsymbol{\varepsilon}\mathbf{E}$.
+When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal structure, the relationship $\mathbf{D^{\prime}}=\boldsymbol{\varepsilon}^{\prime}\mathbf{E^{\prime}}$ holds, where $\mathbf{D^{\prime}}=\mathbf{Q}\mathbf{D}$ and $\mathbf{E^{\prime}}=\mathbf{Q}\mathbf{E}$.
 Substituting these into the equation, we have:
-$\mathbf{Q}\mathbf{D}=\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E}\to\mathbf{D}=\mathbf{Q}^{\top}\bm{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E},$ (21)
-which implies that the dielectric tensor transforms under the $O(3)$ group transformation as: $\bm{\varepsilon}^{\prime}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}$.
+$\mathbf{Q}\mathbf{D}=\boldsymbol{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E}\to\mathbf{D}=\mathbf{Q}^{\top}\boldsymbol{\varepsilon}^{\prime}\mathbf{Q}\mathbf{E},$ (21)
+which implies that the dielectric tensor transforms under the $O(3)$ group transformation as: $\boldsymbol{\varepsilon}^{\prime}=\mathbf{Q}\boldsymbol{\varepsilon}\mathbf{Q}^{\top}$.
 This transformation principle extends similarly to other crystal tensors, as follows.
 Piezoelectric tensor.
-The piezoelectric tensor $\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ to the electric displacement field $\mathbf{D}\in\mathbb{R}^{3}$ within the material.
-Mathematically, this relationship is expressed as $\mathbf{D}_{i}=\sum_{jk}\mathbf{e}_{ijk}\bm{\epsilon}_{jk}$, with $i,j,k\in\{1,2,3\}$.
-When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and electric displacement field is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\mathbf{D}_{i}^{\prime}=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$.
-The relation is then reformulated as $\mathbf{D}_{i}^{\prime}=\sum_{jk}\mathbf{e}_{ijk}^{\prime}\bm{\epsilon}_{jk}^{\prime}$.
-Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$.
+The piezoelectric tensor $\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$ describes the relationship between the applied strain $\boldsymbol{\epsilon}\in\mathbb{R}^{3\times 3}$ to the electric displacement field $\mathbf{D}\in\mathbb{R}^{3}$ within the material.
+Mathematically, this relationship is expressed as $\mathbf{D}_{i}=\sum_{jk}\mathbf{e}_{ijk}\boldsymbol{\epsilon}_{jk}$, with $i,j,k\in\{1,2,3\}$.
+When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and electric displacement field is transformed to $\boldsymbol{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\epsilon}_{mn}$ and $\mathbf{D}_{i}^{\prime}=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$.
+The relation is then reformulated as $\mathbf{D}_{i}^{\prime}=\sum_{jk}\mathbf{e}_{ijk}^{\prime}\boldsymbol{\epsilon}_{jk}^{\prime}$.
+Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\boldsymbol{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\boldsymbol{\epsilon}^{\prime}_{mn}$.
 Consequently, $\mathbf{D}_{i}^{\prime}$ can be represented as
 $\mathbf{D}_{i}^{\prime}$ $=\sum_{\ell}\mathbf{Q}_{i\ell}\mathbf{D}_{\ell}$ (22)
-$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}\bm{\epsilon}_{jk}$
-$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}(\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn})$
-$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{mn}\mathbf{e}_{\ell mn}(\sum_{jk}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}^{\prime}_{jk})\quad(exchange\,sign,m\leftrightarrow j,n\leftrightarrow k)$
-$=\sum_{jk}\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}\bm{\epsilon}_{jk}^{\prime}$
+$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}\boldsymbol{\epsilon}_{jk}$
+$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{jk}\mathbf{e}_{\ell jk}(\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\boldsymbol{\epsilon}^{\prime}_{mn})$
+$=\sum_{\ell}\mathbf{Q}_{i\ell}\sum_{mn}\mathbf{e}_{\ell mn}(\sum_{jk}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\epsilon}^{\prime}_{jk})\quad(exchange\,sign,m\leftrightarrow j,n\leftrightarrow k)$
+$=\sum_{jk}\sum_{lmn}\mathbf{Q}_{il}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\mathbf{e}_{lmn}\boldsymbol{\epsilon}_{jk}^{\prime}$
 Therefore, under the $O(3)$ group transformation $\mathbf{Q}$, the transformed piezoelectric tensor $\mathbf{e}_{ijk}^{\prime}$ is given by:
@@
 Elastic tensor.
-The elastic tensor $C\in\mathbb{R}^{3\times 3\times 3\times 3}$ describes the relationship between the applied strain $\bm{\epsilon}\in\mathbb{R}^{3\times 3}$ and the stress tensor $\sigma\in\mathbb{R}^{3\times 3}$ within the material.
-This relationship is expressed as $\bm{\sigma}_{ij}=\sum_{k\ell}C_{ijk\ell}\bm{\epsilon}_{k\ell}$, with $i,j,k,\ell\in\{1,2,3,4\}$.
-When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and stress tensor is transformed to $\bm{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\epsilon}_{mn}$ and $\bm{\sigma}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\bm{\sigma}_{mn}$.
-Under this transformation, the relationship becomes $\bm{\sigma}_{ij}^{\prime}=\sum_{k\ell}C_{ijk\ell}^{\prime}\bm{\epsilon}_{k\ell}^{\prime}$.
-Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\bm{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\bm{\epsilon}^{\prime}_{mn}$.
-Based on above equations, the transformed stress tensor $\bm{\sigma}_{ij}^{\prime}$ can be expressed as:
-$\bm{\sigma}_{ij}^{\prime}$ $=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\bm{\sigma}_{mn}$ (24)
-$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\bm{\epsilon}_{pq}$
-$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\sum_{k\ell}\mathbf{Q}_{kp}\mathbf{Q}_{\ell q}\bm{\epsilon}^{\prime}_{k\ell}$
-$=\sum_{k\ell}\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}\bm{\epsilon}^{\prime}_{k\ell}$
+The elastic tensor $C\in\mathbb{R}^{3\times 3\times 3\times 3}$ describes the relationship between the applied strain $\boldsymbol{\epsilon}\in\mathbb{R}^{3\times 3}$ and the stress tensor $\sigma\in\mathbb{R}^{3\times 3}$ within the material.
+This relationship is expressed as $\boldsymbol{\sigma}_{ij}=\sum_{k\ell}C_{ijk\ell}\boldsymbol{\epsilon}_{k\ell}$, with $i,j,k,\ell\in\{1,2,3,4\}$.
+When an $O(3)$ group transformation $\mathbf{Q}$ is applied to the crystal, the strain tensor and stress tensor is transformed to $\boldsymbol{\epsilon}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\epsilon}_{mn}$ and $\boldsymbol{\sigma}_{jk}^{\prime}=\sum_{mn}\mathbf{Q}_{jm}\mathbf{Q}_{kn}\boldsymbol{\sigma}_{mn}$.
+Under this transformation, the relationship becomes $\boldsymbol{\sigma}_{ij}^{\prime}=\sum_{k\ell}C_{ijk\ell}^{\prime}\boldsymbol{\epsilon}_{k\ell}^{\prime}$.
+Since $\mathbf{Q}$ is orthogonal matrix ($\mathbf{Q}^{-1}=\mathbf{Q}^{\top}$), we have $\boldsymbol{\epsilon}_{jk}=\sum_{mn}\mathbf{Q}_{mj}\mathbf{Q}_{nk}\boldsymbol{\epsilon}^{\prime}_{mn}$.
+Based on above equations, the transformed stress tensor $\boldsymbol{\sigma}_{ij}^{\prime}$ can be expressed as:
+$\boldsymbol{\sigma}_{ij}^{\prime}$ $=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\boldsymbol{\sigma}_{mn}$ (24)
+$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\boldsymbol{\epsilon}_{pq}$
+$=\sum_{mn}\mathbf{Q}_{im}\mathbf{Q}_{jn}\sum_{pq}C_{mnpq}\sum_{k\ell}\mathbf{Q}_{kp}\mathbf{Q}_{\ell q}\boldsymbol{\epsilon}^{\prime}_{k\ell}$
+$=\sum_{k\ell}\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}\boldsymbol{\epsilon}^{\prime}_{k\ell}$
 Therefore, under the $O(3)$ group transformation $\mathbf{Q}$, $C_{ijkl}^{\prime}$ is represented as:
@@
 When GoeCTP is used to predict elastic tensor, Equation 6 becomes $C_{ijkl}^{\mathrm{final}}=\sum_{mnpq}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq}$.
-A.9 Tensor properties symmetry
+A.5 Tensor properties symmetry
 Tensor properties describe the material’s response to external physical fields (such as electric fields or mechanical stress).
 In materials with symmetry, this response must adhere to the symmetry requirements of the material.
-As demonstrated by Yan et al.
-[51], when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$.
-Therefore, the corresponding tensor properties also remain unchanged, i.e., $\bm{\varepsilon}=\mathbf{R}\bm{\varepsilon}\mathbf{R}^{\top}$.
+As demonstrated by yanspace, when the space group transformation $\mathbf{R}$ is applied to the corresponding crystal structure $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the crystal remains unchanged, i.e., $(\mathbf{A},\mathbf{F},\mathbf{L})=(\mathbf{A},\mathbf{F},\mathbf{R}\mathbf{L})$.
+Therefore, the corresponding tensor properties also remain unchanged, i.e., $\boldsymbol{\varepsilon}=\mathbf{R}\boldsymbol{\varepsilon}\mathbf{R}^{\top}$.
 Thus, crystal symmetry imposes strict constraints on the components of the tensor, leading to the simplification or elimination of many components, reducing the number of independent components.
 Crystal system Number of independent elements Dielectric tensor
-Cubic 1 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
-Tetragonal & Hexagonal & Trigonal 2 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
-Orthorhombic 3 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{22}&0\\ 0&0&\bm{\varepsilon}_{33}\end{pmatrix}$
-Monoclinic 4 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&\bm{\varepsilon}_{13}\\ 0&\bm{\varepsilon}_{22}&0\\ \bm{\varepsilon}_{13}&0&\bm{\varepsilon}_{33}\end{pmatrix}$
-Triclinic 6 $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&\bm{\varepsilon}_{12}&\bm{\varepsilon}_{13}\\ \bm{\varepsilon}_{12}&\bm{\varepsilon}_{22}&\bm{\varepsilon}_{23}\\ \bm{\varepsilon}_{13}&\bm{\varepsilon}_{23}&\bm{\varepsilon}_{33}\end{pmatrix}$
+Cubic 1 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
+Tetragonal & Hexagonal & Trigonal 2 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
+Orthorhombic 3 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{22}&0\\ 0&0&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
+Monoclinic 4 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&\boldsymbol{\varepsilon}_{13}\\ 0&\boldsymbol{\varepsilon}_{22}&0\\ \boldsymbol{\varepsilon}_{13}&0&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
+Triclinic 6 $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&\boldsymbol{\varepsilon}_{12}&\boldsymbol{\varepsilon}_{13}\\ \boldsymbol{\varepsilon}_{12}&\boldsymbol{\varepsilon}_{22}&\boldsymbol{\varepsilon}_{23}\\ \boldsymbol{\varepsilon}_{13}&\boldsymbol{\varepsilon}_{23}&\boldsymbol{\varepsilon}_{33}\end{pmatrix}$
 Table 7: Number of independent components in the dielectric tensor for different crystal systems.
@@
 The $3\times 3$ dielectric tensor has a minimum of 1 and a maximum of 6 independent elements for various types of systems due to the crystal symmetry.
-The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 8 [28].
+The number of independent components in the dielectric tensor for different crystal systems is shown in Tabel 7 (mao2024dielectric).
 Voigt notation for elastic tensor.
@@
 The elastic tensor has a minimum of 3 and a maximum of 21 independent elements for various types of systems.
-Voigt notation is a compact way to represent these independent components of tensor properties [14].
-According to the rules $11\to 1\mathrm{~{};~{}}22\to 2\mathrm{~{};~{}}33\to 3\mathrm{~{};~{}}23,~{}32\to 4\mathrm{~{};~{}}31,~{}13\to 5\mathrm{~{};~{}}12,~{}21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix [46, 37]:
+Voigt notation is a compact way to represent these independent components of tensor properties (itin2013constitutive).
+According to the rules $11\to 1\mathrm{~;~}22\to 2\mathrm{~;~}33\to 3\mathrm{~;~}23,~32\to 4\mathrm{~;~}31,~13\to 5\mathrm{~;~}12,~21\to 6$, the elastic tensor in Voigt notation is a $6\times 6$ symmetric matrix (wen2024equivariant; ran2023velas):
 $C=\begin{pmatrix}C_{1111}&C_{1122}&C_{1133}&C_{1123}&C_{1131}&C_{1112}\\ C_{1122}&C_{2222}&C_{2233}&C_{2223}&C_{2231}&C_{2212}\\ C_{1133}&C_{2233}&C_{3333}&C_{3323}&C_{3331}&C_{3312}\\ C_{1123}&C_{2223}&C_{3323}&C_{2323}&C_{2331}&C_{2312}\\ C_{1131}&C_{2231}&C_{3331}&C_{2331}&C_{3131}&C_{3112}\\ C_{1112}&C_{2212}&C_{3312}&C_{2312}&C_{3112}&C_{1212}\end{pmatrix}\to\begin{pmatrix}C_{11}&C_{12}&C_{13}&C_{14}&C_{15}&C_{16}\\ C_{12}&C_{22}&C_{23}&C_{24}&C_{25}&C_{26}\\ C_{13}&C_{23}&C_{33}&C_{34}&C_{35}&C_{36}\\ C_{14}&C_{24}&C_{34}&C_{44}&C_{45}&C_{46}\\ C_{15}&C_{25}&C_{35}&C_{45}&C_{55}&C_{56}\\ C_{16}&C_{26}&C_{36}&C_{46}&C_{56}&C_{66}\end{pmatrix}$
-The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to Wen et al. [46], Ran et al. [37]).
+The number of independent components in the elastic tensor for different crystal systems is shown in Tabel 8 (partial data shown; for more details, refer to wen2024equivariant; ran2023velas).
 Crystal system Number of independent elements Elastic tensor
@@
 Voigt notation for piezoelectric tensor.
-The number of independent components in the piezoelectric tensor for different crystal systems is shown in Tabel 10 (partial data shown; for more details, refer to De Jong et al.
-[4], Gorfman and Zhang [8]).
+The number of independent components in the piezoelectric tensor for different crystal systems is shown in Tabel 9 (partial data shown; for more details, refer to de2015database; gorfman2024piezoelectric).
 Crystal system point groups Number of independent elements Piezoelectric tensor
@@
 Table 9: Number of independent components in the piezoelectric tensor for different crystal systems.
-A.10 Experimental details
+A.6 Experimental details
 Datasets details.
-The dataset for dielectric tensor and piezoelectric tensor is derived from the data processed by Yan et al.
-[51], sourced from the JARVIS-DFT database.
-Similarly, the dataset for elastic tensor is obtained from the dft_3d data within the JARVIS-DFT database111 https://pages.nist.gov/jarvis/databases/.
+The dataset for dielectric tensor and piezoelectric tensor is derived from the data processed by yanspace, sourced from the JARVIS-DFT database.
+Since nearly half of the crystal samples in the piezoelectric dataset have piezoelectric tensor labels that are entirely zero, training GoeCTP directly on this dataset can lead to severe overfitting, thereby compromising the reliability of the experimental results.
+Therefore, we filtered out all crystal samples with zero-valued piezoelectric tensor labels from piezoelectric dataset.
+For the elastic tensor prediction task, the dataset is also obtained from the JARVIS-DFT database and is publicly accessible as dft_3d through the jarvis-tools package  [^3] https://pages.nist.gov/jarvis/databases/.
 The statistical details of the datasets are presented in Table 10.
@@
 we only predict the $3\times 6$ matrix.
-In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor [14, 46].
+In the elastic tensor dataset, the elastic tensor is represented using Voigt notation as a $6\times 6$ symmetric matrix, rather than a $3\times 3\times 3\times 3$ fourth-order tensor (itin2013constitutive; wen2024equivariant).
 Therefore, we predict the $6\times 6$ matrix during usage (see Appendix A.5 for details).
@@
 Dielectric 4713 14.7 18.2 Unitless
-Piezoelectric 4998 0.43 3.09 $\text{C}/\text{m}^{2}$
+Piezoelectric 2701 0.79 4.03 $\text{C}/\text{m}^{2}$
 Elastic 25110 306.4 238.4 GPa
@@
 The experiments were performed using a single NVIDIA GeForce RTX 3090 GPU.
-For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al.
-[51].
-For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio.
-During model training, Huber loss, AdamW [24], a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
-Hyperparameter settings of GoeCTP.
-When constructing the crystal graph, we used the 16th nearest atom to determine the cutoff radius.
-For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector.
-In the dielectric and piezoelectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector.
-For the prediction module, eComFormer, in the dielectric and piezoelectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer.
-For both dielectric, piezoelectric, and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64.
-In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output.
-For the piezoelectric prediction task, the 18-dimensional vector features output by the prediction module are reconstructed into a $3\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
-For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
-Hyperparameter settings of GMTNet, ETGNN, and MEGNET.
-Following Yan et al.
-[51], we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks.
+For benchmarking, we directly utilized the codebases for ETGNN and GMTNet as provided by yanspace.
+For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio (yanspace).
+[…]
+Hyperparameter settings of GoeCTP (C.
+Fra.).
+For all tensor prediction tasks, the configuration of the prediction module, CrystalFramer (ito2025rethinking), including the input atom-embedding layer, self-attention blocks, and other architectural components, was kept identical to the setup described in the original work (ito2025rethinking) on the JARVIS dataset.
+For the dielectric tensor prediction task, the model was trained using an L1 loss function with a learning rate of 0.002, a batch size of 256, and 200 training epochs, following a polynomial learning rate decay schedule.
+For the piezoelectric tensor prediction task, the model employed an L1 loss function with a learning rate of 0.002, a batch size of 512, and 1000 training epochs, using a warm-up-free inverse square root schedule (huang2020improving; ito2025rethinking).
+For the elastic tensor prediction task, the model adopted the Huber loss function with a learning rate of 0.001, a batch size of 64, and 500 training epochs, along with a polynomial learning rate decay schedule.
+Across all experiments, the AdamW optimizer (loshchilovdecoupled) was used with parameters $(\beta_{1},\beta_{2})=(0.9,0.98)$ and a weight decay of $10^{-5}$.
+Hyperparameter settings of GMTNet and ETGNN.
+Following yanspace, we trained GMTNet and ETGNN for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks.
 The same polynomial learning rate decay scheduler is used in all experiments.
-A.11 Additional results.
-We assessed the performance of GoeCTP when combined with iComFormer [50] and CrystalFormer [42] (denoted as GoeCTP (iCom.) for the combination) on both the dielectric, piezoelectric, and elastic tensor datasets.
-As presented in Table 12, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.).
-As shown in Table 13, similar to the dielectric tensor dataset, GoeCTP (eCom.) achieved the best prediction accuracy on the piezoelectric dataset, but GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.).
-In contrast, as shown in Table 14, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
-MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
-Fnorm $\downarrow$ 3.71 3.40 3.28 3.23 3.40 3.53
-EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2% 81.7% 80.1%
-EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8% 53.8% 52.9%
-EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5% 32.3% 30.6%
-Total Time (s) $\downarrow$ 663 1325 1611 616 535 645
-Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048 0.042 0.202
+For the dielectric tensor prediction task and the piezoelectric tensor prediction task, the network configuration of GMTNet followed that described in the original paper, comprising two layers for node-invariant feature updating and three layers for equivariant message passing.
+For the elastic tensor prediction task, the configuration was adjusted to include two layers for node-invariant feature updating and five layers for equivariant message passing.
+A.7 Additional results.
+We assessed the performance of GoeCTP when combined with iComFormer (yancomplete) and CrystalFormer (taniaicrystalformer) on both the dielectric and elastic tensor datasets.
+We additionally included the baseline method MEGNet (chen2019graph; morita2020modeling) in our comparative experiments.
+Additional comparison results for the dielectric tensor dataset are presented in Table 11, and those for the elastic tensor dataset are presented in Table 12.
+MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.) GoeCTP (C. Fra.)
+Fnorm $\downarrow$ 3.71 3.40 3.28 3.23 3.40 3.53 3.05
+EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2% 81.7% 80.1% 86.4%
+EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8% 53.8% 52.9% 62.6%
+EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5% 32.3% 30.6% 43.5%
+Total Time (s) $\downarrow$ 663 1325 1611 616 535 645 1976
 Table 11: Additional comparison of performance metrics on dielectric dataset.
-MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
-Fnorm $\downarrow$ 0.504 0.515 0.450 0.431 0.448 0.457
-EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9% 45.5% 42.7%
-EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9% 44.9% 42.1%
-EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7% 38.3% 41.5%
-Total Time (s) $\downarrow$ 943 1220 5771 938 909 2787
-Time/batch (s) $\downarrow$ 0.074 0.095 0.45 0.073 0.071 0.218
-Table 13: Additional comparison of performance metrics on piezoelectric dataset.
-MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
-Fnorm $\downarrow$ 143.86 123.64 117.62 107.11 102.80 107.44
-EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5% 46.7% 43.5%
-EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3% 18.6% 15.8%
-EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2% 8.2% 7.9%
-Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422 4035 7891
-Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189 0.315 0.616
+MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.) GoeCTP (C. Fra.)
+Fnorm $\downarrow$ 143.86 123.64 117.62 107.11 102.80 107.44 95.98
+EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5% 46.7% 43.5% 49.7%
+EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3% 18.6% 15.8% 20.3%
+EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2% 8.2% 7.9% 9.8%
+Total Time (s) $\downarrow$ 2899 4448 >> 36000 2422 4035 7891 17197
 Table 12: Additional comparison of performance metrics on elastic dataset.
-A.12 The limitations of our method and its broader impact
-Limitations.
-(1)As discussed in the main text, GoeCTP is a plug-and-play $O(3)$-equivariant framework designed to enhance the backbone network’s ability to achieve equivariant predictions.
-Consequently, the performance of our method is inherently dependent on the capabilities of the backbone network.
-If the backbone lacks sufficient predictive power, the combined framework may not fully surpass the state-of-the-art tensor prediction networks in all evaluation metrics.
-For instance, as shown in Table 12, when our framework is integrated with CrystalFormer [42], although the computational speed significantly exceeds that of GMTNet, the Fnorm, EwT 25%, and EwT 10% values are lower than those achieved by GMTNet.
-(2)As demonstrated in Higham [12] and Proposition 3.5, polar decomposition applied to a $3\times 3$ invertible matrix $\mathbf{L}$ produces a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$.
-For 3D crystal structures, the lattice matrix $\mathbf{L}$ is always full-rank (i.e.
-invertible), therefore, ensuring the applicability of our method to 3D crystal systems.
-However, for certain special cases, such as 2D crystals with single-layer structures [32, 39], the rank of the lattice matrix $\mathbf{L}$ may be less than 3.
-In these scenarios, directly applying polar decomposition may not yield a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$ and a unique canonical form $\mathbf{H}$.
-This limitation could cause our method to fail, as crystals with different spatial orientations cannot be consistently adjusted to a unique canonical form, thereby preventing the method from achieving the desired $O(3)$-equivariance.
-Broader impact.
-This work can be applied to screen existing crystal material databases for potential material candidates, facilitating the discovery of new materials with desirable tensor properties, such as high-dielectric materials.
-Therefore, our work has the potential to make a significant impact on materials science.
-A.13 How to further utilize the tensor properties symmetry
+A.8 How to further utilize the tensor properties symmetry
 In this section, we will use the dielectric tensor as an example to briefly discuss how to further utilize the symmetry of tensor properties.
 Utilizing symmetry constraints for zero elements.
-To demonstrate the role of symmetry constraints, we present an example of the GoeCTP prediction results in Table 15.
+To demonstrate the role of symmetry constraints, we present an example of the GoeCTP (eCom.) prediction results in Table 13.
 Within a reasonable margin of error, the predictions align well with the expected constraints.
 Label Prediction Cubic dielectric tensor
-$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.252&0.016&0.008\\ 0.016&2.230&0.007\\ 0.008&0.007&2.262\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
+$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.252&0.016&0.008\\ 0.016&2.230&0.007\\ 0.008&0.007&2.262\end{pmatrix}$ $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
 Table 13: An example of GoeCTP prediction results
@@
 Therefore, to achieve a higher success rate for zero elements, we added a ReLU activation function to the output layer of the network to improve the success rate (this applies only to cases where tensor elements are greater than or equal to zero).
-The results after retraining GoeCTP are as shown in Table 17, Table 18, and Table 19.
+The results after retraining GoeCTP (eCom.) are as shown in Table 15, Table 16, and Table 17.
 Our method predicts zero elements more accurately while maintaining overall performance.
 Label Prediction Cubic dielectric tensor
-$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.237&0.000&0.000\\ 0.000&2.283&0.000\\ 0.000&0.000&2.228\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
+$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.237&0.000&0.000\\ 0.000&2.283&0.000\\ 0.000&0.000&2.228\end{pmatrix}$ $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
 Table 15: An example of GoeCTP (ReLU) prediction results
@@
 Table 17: Performance comparisons between GoeCTP and GoeCTP (ReLU) on the dielectric dataset.
-This simple modification enables GoeCTP to more accurately predict the zero elements in dielectric tensors caused by the space group constraints.
+This simple modification enables GoeCTP (eCom.) to more accurately predict the zero elements in dielectric tensors caused by the space group constraints.
 Utilizing symmetry constraints for non-zero elements.
@@
 For example, for a cubic crystal, we predict only one independent component and then use the mask from Table 7 to reconstruct it into a $3\times 3$ tensor).
-In Table 15, GoeCTP was modified to directly predict independent components.
+In Table 13, GoeCTP (eCom.) was modified to directly predict independent components.
 The new results are shown in Table 18, which fully satisfy the symmetry constraints.
@@
 Label Prediction Cubic dielectric tensor
-$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.357&0.000&0.000\\ 0.000&2.357&0.000\\ 0.000&0.000&2.357\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
+$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.357&0.000&0.000\\ 0.000&2.357&0.000\\ 0.000&0.000&2.357\end{pmatrix}$ $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
 Table 18: An example of GoeCTP (predicting the independent components) prediction results
@@
 In the second method, once the network training is completed, the network output is weighted using the standard tensor property format, referred to as a mask, as detailed in Table 7.
-The application of this method to the piezoelectric and elastic datasets is demonstrated in Tables 22 and 23.
+The application of this method to the elastic datasets is demonstrated in Table 20.
 The results indicate that incorporating symmetry-based weighting effectively improves the performance of GoeCTP.
-GoeCTP GoeCTP (weighted)
-Fnorm $\downarrow$ 0.431 0.389
-EwT 25% $\uparrow$ 46.9% 48.5%
-EwT 10% $\uparrow$ 42.9% 46.7%
-EwT 5% $\uparrow$ 39.7% 46.1%
-Table 22: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the piezoelectric dataset.
 GoeCTP GoeCTP (weighted)
@@
 Table 20: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the elastic dataset.
-A.14 Analysis of discontinuity in canonicalization
-Previous studies [6] have shown that canonicalization can often introduce discontinuities into the model.
+A.9 Analysis of discontinuity in canonicalization
+Previous studies (dymequivariant) have shown that canonicalization can often introduce discontinuities into the model.
 To assess the impact of such discontinuities on our method, we analyze our proposed method from both theoretical and empirical perspectives.
-From theoretical perspective: The following proposition, presented in Appendix B of Dym et al. [6], characterizes the existence of a continuous canonicalization under orthogonal group actions:
-Proposition A.3.
+From theoretical perspective: The following proposition, presented in Appendix B of dymequivariant, characterizes the existence of a continuous canonicalization under orthogonal group actions:
+Proposition 3.
 Consider the action of the orthogonal group $O(d)$ on $\mathbf{X}\in\mathbb{R}^{n\times d}$.
@@
 Proof.
-According to Proposition A.3, for a matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$, its continuous canonicalization norm is given by $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}$.
+According to Proposition 3, for a matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$, its continuous canonicalization norm is given by $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}$.
 In this work, the polar decomposition is applied to the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H},\mathbf{L}\in\mathbb{R}^{3\times 3}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{H}$ is the positive semi-definite Gram matrix (which serves as the canonicalization norm).
-Based on the properties of positive semi-definite matrices, our method is actually equivalent to the canonicalization approach described in Proposition A.3, i.e.
+Based on the properties of positive semi-definite matrices, our method is actually equivalent to the canonicalization approach described in Proposition 3, i.e.
 $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}=\left((\mathbf{Q}\mathbf{H})^{\top}\mathbf{Q}\mathbf{H}\right)^{1/2}=\left(\mathbf{H}^{\top}\mathbf{H}\right)^{1/2}=\mathbf{H}$.
@@
 Therefore, our canonicalization via polar decomposition is consistent with the continuous canonicalization.
-From empirical perspective: To further demonstrate how this discontinuity affects the accuracy of tensor property predictions, we conducted a set of comparative experiments in which we replaced the polar decomposition in GoeCTP with QR decomposition [22] for evaluation.
+From empirical perspective: To further demonstrate how this discontinuity affects the accuracy of tensor property predictions, we conducted a set of comparative experiments in which we replaced the polar decomposition in GoeCTP (eCom.) with QR decomposition (linequivariance) for evaluation.
 In this QR decomposition case, the lattice matrix is decomposed as $\mathbf{L}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{R}$ is an upper triangular matrix that serves as the canonicalization norm.
-However, $\mathbf{R}$ does not satisfy Proposition A.3, and therefore does not constitute a continuous canonicalization approach.
+However, $\mathbf{R}$ does not satisfy Proposition 3, and therefore does not constitute a continuous canonicalization approach.
 To assess the sensitivity of each approach to small structural perturbations, we used the crystal $CdI_{2}$ (JARVIS_ID: JVASP-29631) as the model input, and then perturbed its lattice matrix (specifically, scaling down the first row of the lattice matrix) to observe the relative changes in the output (specifically the variations in the Fabenius norm of the output tensor).
@@
 This highlights the robustness of the continuous canonicalization afforded by polar decomposition, demonstrating its superiority in maintaining prediction stability and structural consistency under small input variations.
-A.15 How to design other canonical forms
+A.10 How to design other canonical forms
 In fact, the canonical form obtained via polar decomposition provides a foundation from which infinitely many alternative canonical forms can be derived.
@@
 This formulation highlights the inherent flexibility in defining canonical forms.
-A.16 Impact of different canonical form on tensor property prediction
+A.11 Impact of different canonical form on tensor property prediction
 As discussed in Appendix A.10, there exist infinitely many possible canonical forms.
-To investigate the impact of different canonical forms on tensor property prediction in crystalline materials, we first present the results comparing the performance of GoeCTP with QR decomposition [22] and GoeCTP without canonicalization.
+To investigate the impact of different canonical forms on tensor property prediction in crystalline materials, we first present the results comparing the performance of GoeCTP (eCom.) with QR decomposition (linequivariance) and GoeCTP (eCom.) without canonicalization.
 The corresponding results are summarized in the table below:
@@
 One possible explanation is that canonical forms typically exhibit structured representations, such as the symmetric positive semi-definite matrix obtained via polar decomposition or the upper triangular matrix from QR decomposition, which may facilitate the model’s ability to capture relevant features.
-Furthermore, the use of a canonical form aligns all atoms within a crystal structure to a unified global coordinate system, thereby enabling equivariant tensor prediction.
-However, recent studies [43, 23, 15] have demonstrated that employing adaptive local frames, i.e.
-constructed based on individual atomic environments, can further enhance the accuracy of property predictions.
-Motivated by these insights, we extend the GoeCTP framework by incorporating local frames [43] (also referred to as local canonicalization).
-It is important to emphasize that GoeCTP can be seamlessly integrated with any prediction network, including those based on local frames, while preserving its equivariant prediction capability.
-The corresponding results are presented below.
-GoeCTP (local frame+polar) GoeCTP (polar)
-Fnorm $\downarrow$ 3.11 3.23
-EwT 25% $\uparrow$ 83.9% 83.2%
-EwT 10% $\uparrow$ 58.4% 56.8%
-EwT 5% $\uparrow$ 33.5% 35.5%
-Total Time (s) $\downarrow$ 1979 616
-Table 26: Predictive performance comparisons between GoeCTP (local frame+polar) and GoeCTP (polar) on the dielectric dataset.
-The results indicate that incorporating local frames can enhance predictive performance. However, this improvement comes at the expense of increased computational cost.